### v1

# Selective Variable Convolution Meets Dynamic Content Guided Attention for Infrared Small Target Detection

Abstract
Infrared Small Target Detection (IRSTD) system aims to identify small targets in complex backgrounds. Due to the convolution operation in Convolutional Neural Networks (CNNs), applying traditional CNNs to IRSTD presents challenges, since the feature extraction of small targets is often insufficient, resulting in the loss of critical features. To address these issues, we propose a dynamic content guided attention multiscale feature aggregation network (DCGANet), which adheres to the attention principle of ’coarse-to-fine’ and achieves high detection accuracy. First, we propose a selective variable convolution (SVC) module that integrates the benefits of standard convolution, irregular deformable convolution, and multi-rate dilated convolution. This module is designed to expand the receptive field and enhance non-local features, thereby effectively improving the discrimination of targets from backgrounds. Second, the core component of DCGANet is a two-stage content guided attention module. This module employs two-stage attention mechanism to initially direct the network’s focus to salient regions within the feature maps and subsequently determine whether these regions correspond to targets or background interference. By retaining the most significant responses, this mechanism effectively suppresses false alarms. Additionally, we propose adaptive dynamic feature fusion (ADFF) module to substitute for static feature cascading. This dynamic feature fusion strategy enables DCGANet to adaptively integrate contextual features, thereby enhancing its ability to discriminate true targets from false alarms. DCGANet has achieved new benchmarks across multiple datasets.
Index Terms:
Infrared small target, selective variable convolution, dynamic content guided attention, adaptive dynamic feature fusion.
I Introduction
Infrared imaging technology, a crucial component of infrared monitoring systems, excels at capturing thermal signatures of targets and demonstrates robust performance in low-light and foggy environments. In military [1], ocean rescue [2], and environmental monitoring [3] applications, targets of interest at long imaging distances often appear as small spots [3]. Consequently, infrared small target detection (IRSTD) is essential for extracting valuable information from complex backgrounds.
Figure 1: The complex background faced by image target detection, the red and green bounding boxes are zoom in area: (a)Dense target of different types such as pedestrians and cars, (b)Target occlusion by other objects such as trees, (c)Strong lighting conditions of sunlight, and (d) Dim lighting conditions during nights.
Despite the critical role of IRSTD in infrared warning systems, as illustrated in Fig. 1, two fundamental challenges remain in the feature extraction of infrared small targets. First, the low resolution of infrared imaging systems results in a paucity of texture and color information. Second, the inherent limitations of imaging small targets occupy an exceedingly small number of pixels. These challenges substantially impede practical performance in complex environments [4]. The primary issues impeding the detection of infrared small targets are as follows:
  1. 1.
    Insufficient features: Infrared small targets often lack distinct structural features due to their low resolution and the absence of texture and color information, making it difficult to differentiate them from complex backgrounds.
  2. 2.
    Low signal-to-noise ratio: The weak infrared radiation signal from small targets, combined with low contrast against background noise, increases the difficulty of detection and recognition.
Addressing these challenges is essential for achieving robust and precise detection in infrared warning systems. It is imperative to suppress background clutter by enhancing the extraction of inherent target features and integrating contextual analysis, which is pivotal to improving IRSTD performance.
Traditional IRSTD methodologies can be categorized into three primary approaches. They are Background Suppression-based (BS-based), Human Visual System-based (HVS-based), and Low Rank Matrix-based (LRM-based) methods. The BS-based methods [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] utilize specialized filters to extract small targets. The HVS-based methods [16, 17, 18, 19, 20, 21, 22, 23] identify small targets by exploiting local information features. The LRM-based methods [24, 25, 26, 27] detect small targets through the decomposition and subsequent reconstruction of infrared images into sparse matrices. However, these traditional methods are limited in their ability to learn from datasets or dynamically adjust parameters in response to changing environmental conditions, leading to suboptimal performance in complex scenarios.
Recent advances in deep learning have yielded substantial improvements for IRSTD. Previous research in deep learning methods emphasized weak feature expressions, such as gradient edge [28, 29, 30, 31], contextual information [32, 33, 34, 35, 36], and local contrast [37, 38, 39]. Despite their progress, fundamental challenges in IRSTD remain unaddressed. Specifically, the current limitations are primarily evident in three aspects:
  1. 1.
    Insufficient Feature Extraction: Current convolution methods have shortcomings in capturing details of infrared small targets. Shallow networks have limited receptive fields, which struggle to capture non-local relationship between small targets and background. Meanwhile, deep networks may lose features due to downsampling, leading to inaccurate discrimination between targets and false alarms, thereby affecting the accuracy of detection.
  2. 2.
    Static Attention for Interference Suppression: Traditional convolution and pooling strategies often struggle to effectively separate background and target features, while the saliency of infrared targets may be distributed in specific channels (such as thermal radiation intensity, gradient changes, etc.). However, existing attention mechanisms (such as SE [10], CBAM [34]) model channel relationships statically or at a single scale, causing target information to be masked by background signals, reducing detection sensitivity.
  3. 3.
    Static Feature Fusion: Current feature fusion methods, such as skip connected feature addition or concatenation [40], ignore the differences in various target features at different levels, which can lead to shallow-level noise contamination of deep semantics features, and thus limit the network’s adaptability in dynamic scenes. This further leads to a high false positive rate, affecting the reliability of the detection.
These limitations collectively impede robust detection in real-world scenarios, where targets exhibit dynamic scales and unpredictable noise patterns.
Facing the challenge of IRSTD, existing convolutional neural networks often struggle to effectively address the feature extraction of infrared small targets. To overcome these challenges, there is an imperative need for a novel framework, which is derived from three pivotal observations: (1) Infrared small targets require hierarchical dynamic receptive fields to jointly capture edge sharpness and thermal radiation context; (2) Feature fusion should be selectively guided by high-level semantics to suppress noise propagation; (3) Attention mechanisms must incorporate infrared physical priors and consider input channels according to wavelength sensitivity, while adapting to environmental dynamics. Therefore, our framework should not only effectively extract subtle features of infrared small targets through advanced feature extraction techniques but also dynamically optimize multi-stage feature fusion processes. This framework requires a deep understanding of the background environment, rather than simply relying on local image information near the target.
To address the aforementioned challenges, we propose DGCANet, which consists of three core innovations. First, to improve the recognition of real targets, suppress background noise, and discriminate between them and false alarms, we proposed selective variable convolution (SVC), which consists of multi-branch adaptive kernels. The SVC resolving the local-global trade-off, which can accurately extract features of weak targets and significantly reduce background noise. Second, a dynamic content guided attention(DCGA) mechanism was proposed. The DGCA first performs local feature extraction and expanded receptive field extraction using parallel ordinary convolution and dilated convolution. To generate multi-scale coarse-grained attention maps, through channel grouping refinement strategy, the input feature channels are grouped according to physical meanings (such as thermal radiation intensity and gradient texture), and captures insensitive features of weak targets through fine attention allocation. Last, we proposed Adaptive Dynamic Feature Fusion (ADFF) to simulate the attention regulation mechanism of the human vision systems(HVS). The integration of ADFF with a wide receptive field significantly improves the distinguish between real targets and false positives.
The main contributions of this paper are listed as follows:
  1. 1.
    We propose DCGANet, a novel framework that unifies feature learning, dynamic content guided attention, and feature fusion in a hierarchical dynamic pipeline. This framework bridges the gap between rigid CNN designs and the dynamic, sparse infrared target, achieving joint optimization of accuracy, robustness, and interpretability.
  2. 2.
    The core of DCGA is to simulate the process of "rapid scanning to fine focusing" in HVS. By using the physical characteristics of infrared imaging, such as the differences in response to targets in different bands, potential target areas were prioritized and located through channel grouping refinement, while irrelevant backgrounds were suppressed.
  3. 3.
    Our SVC combines deformable convolution, multi-dilated convolution, and standard convolution. This architecture enhances the discrimination between real targets and false alarms, and suppresses background noise through its broad receiving domain.
  4. 4.
    The ADFF substantially enhances the main network’s receptive field capacity to differentiate between actual targets and false positives via its sophisticated dynamic fusion technology.
II Related Work
II-A Infrared Small Target Detection
The methodological evolution of infrared small target detection (IRSTD) is divided into model-driven and data-driven paradigms.
II-A1 Model-Driven Infrared Small Target Detection
Model-driven methods formalize IRSTD as anomaly recognition against backgrounds, leveraging three computational paradigms: filters-based, human visual system (HVS) simulations-based, and sparse matrix reconstruction. The filter-based method [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] mainly focuses on suppressing background clutter and highlighting targets while suppressing isolated noise through filter stacking. For instance, multi-scale Top-Hat [8, 9] filtering algorithm, which combines multi-scale transformation, can improve detection performance. However, when the infrared image contains complex backgrounds, the algorithm will exhibit a lower signal-to-noise ratio (SNR). To further improve the performance, the Human Visual System-Based (HVS-Based) methods [16, 17, 18, 19, 20, 21, 22, 23, 41] utilize the principle of human visual contrast to suppress background brightness and enhance local information of infrared small targets, such as LCM [16], and TLLCM [41]. However, HVS-based methods still struggle to maintain detection stability when faced with high interference from bright edges or points. From another perspective, sparse matrix reconstruction treats small targets as sparse matrices and backgrounds as low-rank matrices [24, 25, 26, 27, 42], detecting small targets by decomposition low-rank matrices. For instance, the Infrared Patch Image (IPI) [42] model divides infrared images into blocks and recombines them into low-rank sparse matrices. Meanwhile, the reweighted infrared patch tensor (RIPT) [24] was proposed to construct low-rank sparse matrices more stably and faster through the alternating direction method. Overall, although these three categories of methods can achieve good detection results for certain images, their performance dramatically decreases when facing complex backgrounds.
II-A2 Data-Driven Infrared Small Target Detection
Data-driven methods, empowered by deep learning advancements, have emerged as a research frontier in IRSTD. Data-driven methods improve original convolutional neural networks (CNNs) with the prior information of infrared small targets. On the one hand, for the enhancement of local information, the attention local contrast (ALCNet) Net [37] was designed to enhance the local contrast of feature maps. Meanwhile, the dynamic attention transformer (DATNet) [43] was proposed to improve the performance through contextual connectivity. Furthermore, the gated-trans unet (GSTUnet) [29] focuses on the local edge of the small target to improve detection performance. On the other hand, in optimizing network architectures for IRSTD, the dense nested attention (DNANet) [34] employs U-shaped network topology with dense skip connections to achieve feature fusion and target enhancement. Additionally, the background-aware feature exchange net (BAFENet) [44] combines background semantics with feature exchange of targets and backgrounds to achieve high-precision detection.
Although deep learning methods have made substantial progress, they still face challenges such as over-parameterized architectures and inflexible receptive fields [43], impending comprehensive capture of the details of small targets. Excessive reliance on local information [31] also limits their ability to model global contextual features, leading to high false alarm rates in noisy scenarios. DCGANet distinguishes itself through two core innovations:
  1. 1.
    Multi-convolutional spatiotemporal feature extraction: Overcoming CNN rigidity via adaptive kernel configurations across spectral-temporal domains.
  2. 2.
    HVS-mimetic dynamic attention: The DCGA module implements ’coase-to-fine’ saliency mapping, mirroring human visual system search patterns through channel-specific attention refinement.
II-B Vision Attention Mechanisms on IRSTD
Vision attention mechanism in IRSTD enhance target feature expression representation by dynamically prioritizing regions, simultaneously amplifying discriminative target and effectively suppressing noise and false alarms. For instance, the local patch network with global attention net (LPNet) [45] improves the detection accuracy of small targets by combining global and local features. The infrared shape network (ISNet) [31] used Taylor finite difference heuristic blocks and bidirectional attention aggregation modules. In addition, the dynamic gradient attention transformer net (DATransNet) [43] further improves detection performance by introducing a transformer detection module and gradient attention.
Despite these advancements, existing methods still lack robust global feature understanding in IRSTD. In contrast, dynamic content guided attention (DCGA) implements a coarse-to-fine attention cascade: it first generates a rough spatial attention map, then refines it into a fine-grained map by integrating spatial and channel attention. This process refines initial spatial attention maps (SIM) for each channel to focus on the most representative features, enhancing the recognition of suspected small targets. In feature fusion, DCGA adopts a shallow-deep feature fusion strategy to improve information flow between network layers, preserving features and enabling effective gradient backpropagation. This dual optimization not only boosts small target detection accuracy but also strengthens false alarm suppression capabilities.
III Method
III-A Overall Architecture
Figure 2: Overall architecture of the proposed DCGANet, built upon a U-Net-like backbone and integrating the key Selective Variable Convolution (SVC), Dynamic Content Guided Attention (DCGA), and Adaptive Dynamic Feature Fusion (ADFF) modules.
The DCGANet architecture, depicted in Fig. 2, is constructed upon a U-Net-like framework and comprises three components: selective variable convolution (SVC), dynamic content guided attention (DCGA), and adaptive dynamic feature fusion (ADFF) modules. Specifically, the encoder path consists of three cascaded layers integrating SVC and DCGA modules in sequence. The SVC replaces standard convolutions to enhance subtle target signatures while expanding receptive fields for augmented context modeling. The DCGA mechanism generates channel-specific spatial importance maps (SIMs) through input-guided attention calibration, enabling per-channel saliency emphasis critical for sub-pixel infrared targets. On the decoder path, two ADFF modules dynamically merge multi-scale encoder features, replacing the original skip-connections and achieving context-aware noise suppression through three-stage fusion.
III-B Selective Variable Convolution
The limited receptive field of traditional convolutional layers in CNNs hinders their ability in IRSTD. To address this, we propose the selective variable convolution (SVC) module, a robust basic block aimed at significantly enhancing the network’s feature learning ability.
As shown in Fig. 3, the SVC combines synergy of the tri-branch features. The first branch employs standard convolution(SConv) to preserve local spatial feature extraction, retaining elementary patterns of small targets and maintaining the stability of fundamental detection performance. The second branch utilizes deformable convolution (DConv), which has been proven effective in edge delineation for irregularly shaped targets, enhancing target features across diverse scales. The third branch, multi-dilation convolution (MDConv), utilizes multi-rate dilated kernels (rates = 2, 4, 8) to extend the receptive field, thereby aggregating multi-scale contexts to suppress structured clutter. The SVC enables accurate extraction of weak target features and significant reduction of background noise.
$X_{s}$ $=\operatorname{SConv}(X_{\text{in}}),$ (1)
$X_{c}$ $=\operatorname{DConv}(X_{\text{in}}),$ (2)
$X_{m}$ $=\operatorname{MDConv}(X_{\text{in}}).$ (3)
where $X_{\text{in}}$ is the input feature, $X_{s}$, $X_{c}$, and $X_{m}$ are the outputs of SConv, DConv, and MDConv, respectively.
The output of these branches is fused through channel-wise concatenation and a $1\times 1$ convolution, yielding consolidated representations enriched in both local precision and global discriminability.
$X_{\text{out}}=\operatorname{Conv}_{1\times 1}\left(\operatorname{Concat}\left(X_{s},X_{c},X_{m}\right)\right),$ (4)
where $X_{\text{out}}$ is the output of SVC module, $X_{s}$, $X_{c}$, and $X_{m}$ are the outputs of SConv, DConv, and MDConv, respectively.
By enhancing the extraction of edges and details through a multi-convolution strategy, SVC effectively boosts recognition capability in IRSTD. The expanded receptive field and deformable feature extraction offer substantial advantages in contextual perception and noise suppression, thereby improving the clarity and reliability of feature extraction.
Figure 3: The structure of the SVC: a parallel branching architecture to perform different types of convolution operations on the input feature maps. This design enables the encoding phase to extract features and preserve details at various scales.
III-C Dynamic Content Guided Attention
While the SVC module expands the receptive field and enhances non-local feature extraction, it still struggles to identify the target when dealing with background clutter. Previous attention mechanisms, such as the static channel-wise or spatial weight separation mechanisms of SE [10], CSAM [34], face two key issues. First, there is a lack of information interaction between channel-wise and spatial attention. Second, single-channel spatial weights cannot adapt to the diverse distribution of multiple feature channels and pose a risk of information loss due to the dilution of key target features.
To address these challenges, we propose the dynamic content guided attention(DCGA) module, which uses a coarse-to-fine approach to generate self-attention maps(SIM) for specific channels, effectively simulating the ability of HVS to focus on highlighting features and suppress false alarms. The DCGA is shown in Fig. 4. The first stage includes coarse-grained attention with interaction between channel-wise and spatial attention, while the second stage includes fine-grained calibration with dynamic SIM generation, obtaining a unique SIM for each single-channel input features in a coarse-to-fine manner. We first calculate the corresponding Wc and Ws by following.
III-C1 Stage I: Coarse-Grained Attention with Joint Channel-Space Fusion
The stage include Channel attention, Spatial Attention, and joint attention.
Channel attention: A dual-bottleneck structure extracts global channel responses by aggregating spatial information through global average pooling:
$W_{c}=\operatorname{\mathcal{C}}_{1\times 1}\left(\operatorname{ReLU}\left(\operatorname{\mathcal{C}}_{1\times 1}\left(X_{GAP}^{\text{c}}\right)\right)\right)$ (5)
where $X_{GAP}^{\text{c}}$ denotes channel-wise squeezed features from global average pooling.
Spatial Attention: Spatial saliency is captured by concatenating multi-perspective features from global average pooling (GAP) and global max pooling (GMP):
$W_{s}=\operatorname{\mathcal{C}}_{7\times 7}\left(\operatorname{Concat}\left(X_{GAP}^{\text{s}},X_{GMP}^{\text{s}}\right)\right).$ (6)
where $X_{GAP}^{\text{s}},X_{GMP}^{\text{s}}$ represent spatial descriptors from global average pooling (GAP) and global max pooling (GMP).
Interaction Attention: we fuse $W_{c}$ and $W_{s}$ together using an element-wise addition operation following broadcast rules to obtain a rough SIM $W_{\text{coa}}$.
$W_{\text{coa}}=W_{c}\oplus W_{s},$ (7)
where $W_{\text{coa}}\in\mathbb{R}^{C\times H\times W}$ is rough SIM, $\oplus$ denotes element-wise addition with automatic dimension expansion.
Figure 4: The proposed structure of the DCGA: employs a parallel approach to equally compute the weights of spatial and channel information. It also incorporates random mixing of channel and spatial information to enhance the capability for target segmentation.
III-C2 Stage II: Fine-Grained Calibration with Dynamic SIM Generation
The fine-grained calibration include channel shuffling fusion and dynamic weight generation.
Channel Shuffling Fusion: The raw feature $X\in\mathbb{R}^{C\times H\times W}$ and coarse attention map $W_{coa}$ are concatenated and then shuffled along the channel dimension [46], establishing inter-channel correlations while preserving spatial details.
Dynamic Weight Generation: A gated convolution layer $\mathcal{GC}_{7\times 7}$ followed by sigmoid activation generates channel-specific SIM :
$W=\sigma\left(\operatorname{\mathcal{GC}}_{7\times 7}\left(\operatorname{CS}\left(\operatorname{Concat}\left(X,W_{\text{coa}}\right)\right)\right)\right)$ (8)
where $\operatorname{CS}(\cdot)$ denotes the channel shuffling operation. This enables content-adaptive calibration by allowing each channel’s SIM to dynamically adjust based on local feature patterns.
DCGA assigns a unique SIM to each channel, guiding the model to focus on the important areas of each channel. Therefore, more useful information encoded in the features can be emphasized to effectively improve IRSTD.
III-D Adaptive Dynamic Feature Fusion Module
The proposed Adaptive Dynamic Feature Fusion(ADFF) module aims to enhance the network’s ability to distinguish between actual targets and similar false alarms. Specifically, we introduced a dynamic space selection mechanism within the original skip connections, as shown in Fig. 5. This mechanism leverages multi-scale features to enable the network to focus on the areas with the richest spatial contextual information, thereby increasing sensitivity to the target area. The ADFF module is specifically designed to process the feature maps of the encoder-decoder pair in stage i, where the feature maps are represented as:
$X_{i},X^{\prime}_{i}\in\mathbb{R}^{C_{i}\times\frac{H}{2^{i}}\times\frac{W}{2^{i}}}$ (9)
where $C_{i}$ denotes channel number, $\frac{H}{2^{i}}$ and $\frac{W}{2^{i}}$ are height and width of feature maps.
Then we applied channel-wise average pooling and max pooling operations to feature maps $X_{i}$and $X^{\prime}_{i}$, represented as $P_{avg}(\cdot)$and $P_{max}(\cdot)$, respectively.
$AD_{i}^{\text{avg}}$ $=\operatorname{AvgPool}\left(\operatorname{Concat}\left[X_{i},X^{\prime}_{i}\right]\right),$ (10)
$AD_{i}^{\text{max}}$ $=\operatorname{MaxPool}\left(\operatorname{Concat}\left[X_{i},X^{\prime}_{i}\right]\right).$ (11)
where $AvgPool$ is average pooling operation, $MaxPool$ is max pooling operation.
These two different pooling operations extract the average and maximum feature descriptors of the space, and the generated spatial pooling features are aggregated into a matrix to facilitate information exchange between different spatial descriptors. Next, we use the convolutional layer $F_{2\to N}(\cdot)$to transform these aggregated features into $N$spatial attention maps. These spatial attention maps can effectively guide the subsequent feature fusion process, highlight salient regions, and suppress redundant background information.
Figure 5: The proposed structure of ADFF: integrates the encoding and decoding stages of the network through a spatially adaptive selection mechanism. It utilizes both maximum and average pooling across channels to achieve precise target extraction.
$AD_{\text{fuse}}=\sigma\left(\operatorname{Conv}_{k\times k}\left(\operatorname{Concat}\left(AD_{i}^{\text{avg}},AD_{i}^{\text{max}}\right)\right)\right),$ (12)
where $\sigma$ denotes the sigmoid function, $\operatorname{Concat}(\cdot)$ is the channel-wise concatenation, and $\operatorname{Conv}_{k\times k}(\cdot)$ represents the convolutional layer (e.g., with a $7\times 7$ kernel, typically outputting a single channel) used to generate the attention map.
The final output $Y_{i}$ of the ADFF module is obtained by first applying the generated attention map $AD_{\text{fuse}}$ to weight the input features $X_{i}$ and $X^{\prime}_{i}$ via element-wise multiplication, and then fusing the results through element-wise addition. This process is formulated as:
$Y_{i}=X_{i}\odot AD_{\text{fuse}}+X^{\prime}_{i}\odot AD_{\text{fuse}}$ (13)
where $\odot$ denotes element-wise multiplication, $Y_{i}$ is the output of ADFF.
The fusion process generates dynamic fusion weights, which can adaptively adjust the importance of each feature map in the final decision. Through this adaptive selective fusion strategy, the ADFF module can automatically identify and integrate the most discriminative features, thereby improving the network’s representation ability of IRSTD.
IV Experiment
[…]
table
TABLE I: Configurations for all comparative experiments
Methods(Source & Year) Key parameters configurations
Background Suppression Methods
NWMTH [12]PR’2010 $R_{o}=9,R_{i}=4$
FKRW [47]TGRS’2019 Windows size: 11, $K=4,p=6,\beta=200$
HVS-based Methods
MPCM [17]PR’2016 $N=1,3,\cdots,9$, threshold $k=3$
WLDM [48]TGRS’2016 $L=4,m=2,n=2$, threshold $k=2$
RLCM [49]GRSL’2018 $k_{1}=[2,5,9],k_{2}=[4,9,16]$, scale: 3, threshold $k=1$
ILCM [50]GRSL’2020 Cell size: $3\times 3$, threshold $k=3$
TLLCM [41]GRSL’2020 Scales:[5,7,9] $\lambda=0.5$, threshold $k=3$
GSWLCM [18]GRSL’2022 Local Window Structures: $[3,5,7,9]$, $\delta=0.01,k=20$
Low-rank and Sparse Decomposition
IPI [42]TIP’2013 Patch size: 50, sliding step: 10, $\lambda=1/\sqrt{\mathrm{min}(m,n)}$, $\epsilon=10^{-7}$
RIPT [24]JSTARS’2017 Patch size: 50, sliding step: 10, $\lambda=2/\sqrt{\mathrm{min}(m,n)}$, $\epsilon=10^{-2}$, $\omega=10^{-7}$
PSTNN [26]RS’2019 Patch size: 40, sliding step: 40, $\lambda=0.6/\sqrt{\mathrm{max}(n_{1},n_{2})\times n_{3}}$, $\epsilon=10^{-7}$
Deep Learning Methods
ACM [32]WACV’2021 Backbone: FPN, layer blocks: [4, 4, 4], channels: [8, 16, 32, 64]
AGPCNet [38]TAES’2023 Backbone: resnet34, scales: [10, 6, 5, 3], reduce ratios: [16, 4], gca type: patch, gca att: post
DNANet [34]TIP’2022 Backbone: resnet18, layer blocks: [2, 2, 2, 2], filter: [16, 32, 64, 128, 256]
ISNet [31]CVPR’2022 Backbone: resnet18, layer blocks: [4, 4, 4], channels: [8, 16, 32, 64], fuse mode: AsymBi
UIUNet [33]TIP’2023 Channels: [64, 128, 256, 512], fuse mode: AsymBi
MTUNet [51]TGRS’2023 Backbone: resnet18, layer blocks: [2, 2, 2, 2], filter: [16, 32, 64, 128, 256]
SeRankDet TGRS’2024 Channels:[64, 128, 256, 512, 1024],o=3
SCTransNet [52] TGRS’2024 Backbone:Swin-Tiny, Layer Num=[2,6,6,2], Head Num=[3,6,12,24]
$\star$ DCGANet (Ours) Channels: [64, 128, 256, 512, 1024]
IV-A Experimental Setup
IV-A1 Datasets
To rigorously evaluate the performance and generalization capabilities of our method, we conduct comprehensive benchmark experiments on three widely recognized infrared small target detection datasets: NUAA-SIRST [32], IRSTD-1K [31], and NUDT-SIRST [34]. These datasets demonstrate considerable diversity in target characteristics, background clutter, and imaging conditions, thereby enabling robust validation of detection methodologies within real-world operational constraints.
  1. 1.
    NUAA-SIRST: The real long-range infrared single-frame dataset comprises 427 images capturing diverse realistic scenes, such as clouds, cities, and oceans. However, the dataset is relatively small in scale, with only 480 target instances, which may limit its effectiveness for training and evaluating models in scenarios that demand high data diversity.
  2. 2.
    IRSTD-1K: IRSTD-1K is designed to address the challenges of multi-scene cross-domain detection, comprising 1,001 highly diverse images. The complexity of its scene types and target shapes imposes significantly higher demands on the model’s generalization capabilities compared with the other two datasets.
  3. 3.
    NUDT-SIRST: The scene-enhanced dataset, constructed from synthetic data, comprises 1,327 images with complex backgrounds, including cities, fields, bright regions, oceans, and clouds. By employing adaptive target size and intensity blurring functions, it generates diverse targets in the form of point-like, patchy, and extended shapes, yielding distribution characteristics that closely resemble those of real-world datasets.
IV-A2 Evaluation Metrics
To quantitative analyze our DCGANet and other methods, we adopt four standard metrics to evaluate detection performance: Intersection over Union (IoU), normalized IoU (nIoU), Probability of detection (Pd), and False alarm rate (Fa). The detailed description of these metrics are as follows:
  1. 1.
    IoU measures the accuracy of object localization by comparing the overlap between the predicted bounding box and the ground truth. It is defined as the ratio of the intersection area to the union area of the two bounding boxes.
    $\text{IoU}=\frac{TP}{T+P-TP}$ (14)
  2. 2.
    nIoU is a standardized adaptation of the IoU metric.
    $\text{nIoU}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP(i)}{T(i)+P(i)-TP(i)}$ (15)
  3. 3.
    Pd quantifies the likelihood of correctly identifying true targets.
    $\text{Pd}=\frac{TP}{TP+FN}$ (16)
  4. 4.
    Fa measures the frequency of incorrect target detections.
    $\text{Fa}=\frac{FP}{FP+TN}$ (17)
where $N$ is the total number of samples, $T$ and $P$ correspond to the number of ground truth and predicted positive pixels, respectively. Additionally, $TP$, $FP$, and $FN$ denote the counts of true positive, false positive, and false negative pixels, respectively.
In addition, the ROC curve illustrates the variation trend of the Probability of detection (Pd) across different False alarm rate (Fa).
IV-A3 Implementation Details
We adopt U-Net as the baseline model, Soft-IoU Loss [53] as the loss function, and AdamW [54] as the optimizer, and incorporating a polynomial learning rate decay strategy to optimize the training process. Additionally, we employ a deep supervision strategy to further enhance model performance. All experiments are conducted using distributed training on a system equipped with four RTX 3090 GPUs. To ensure comprehensive and reliable evaluation, we train independent models on each of the three datasets. The detailed hyperparameter configurations for all comparative methods are summarized in Tab. I. To address the significant differences in dataset distribution and resolution, we customize the hyperparameters for each dataset, with specific configurations provided in Tab. II.
TABLE II: Customized Hyperparameters for Each Dataset
Dataset Epochs Lr Batch Resolution
NUAA-SIRST 500 0.0001 4 512$\times$512
IRSTD-1K 500 0.0001 4 512$\times$512
NUDT-SIRST 500 0.0001 16 256$\times$256
TABLE III: Comparison with Other State-of-the-art methods on four datasets. The $\uparrow$ indicates that the higher the indicator, the better, and the $\downarrow$ indicates that the lower the indicator, the better. The best performance is highlighted in bold, and the second-best is underlined.
Method NUAA-SIRST(Tr=60%) IRSTD-1K(Tr=60%) NUDT-SIRST(Tr=60%)
IoU $\uparrow$ nIoU $\uparrow$ $P_{d}$uparrow $F_{a}$downarrow IoU $\uparrow$ nIoU $\uparrow$ $P_{d}$uparrow $F_{a}$downarrow IoU $\uparrow$ nIoU $\uparrow$ $P_{d}$uparrow $F_{a}$downarrow
                      Background Suppression Methods
NWMTH [12] 15.35 17.00 71.10 55.21 18.31 16.74 49.97 21.06 18.87 16.22 62.84 32.45
FKRW [47] 22.10 28.81 81.41 16.67 10.12 16.83 69.88 24.19 10.04 16.76 72.31 66.90
HVS-based Methods
MPCM [17] 25.64 27.63 68.01 45.07 20.72 23.26 58.78 31.10 27.40 38.95 55.54 11.82
WLDM [48] 5.55 3.37 64.97 21.22 8.16 10.32 54.85 14.79 8.19 11.63 16.30 64.14
RLCM [49] 25.56 27.46 86.34 62.16 16.46 21.83 70.86 67.70 8.21 11.42 70.22 31.67
ILCM [50] 18.42 19.16 38.00 40.63 13.67 11.34 43.32 47.67 5.16 6.48 26.27 88.16
TLLCM [41] 18.20 27.62 82.23 16.18 10.86 18.01 68.95 24.99 13.44 16.98 77.02 69.66
GSWLCM [18] 15.67 13.82 72.48 21.96 12.81 13.46 70.57 13.82 21.08 19.92 69.06 42.78
Low-rank and Sparse Decomposition
IPI [42] 26.67 32.57 83.55 11.97 27.92 30.12 81.27 26.18 25.16 34.64 76.26 43.41
RIPT [24] 11.85 14.91 79.48 22.67 14.71 17.43 79.55 28.35 24.13 33.98 78.54 56.24
PSTNN [26] 22.72 23.50 77.36 29.99 28.83 29.09 82.68 39.48 21.26 24.22 72.78 61.26
Deep Learning Methods
ACM [32] 73.38 74.66 96.06 9.94 60.38 59.83 91.82 21.49 73.86 72.97 97.91 76.22
AGPCNet [38] 77.86 75.23 98.92 34.95 68.30 66.18 94.29 15.55 88.55 88.08 97.40 34.09
DNANet [34] 78.82 79.72 98.63 13.37 69.96 69.35 92.02 23.19 91.32 90.10 97.55 30.71
ISNet[31] 80.83 78.55 98.78 4.93 68.74 61.98 95.80 14.97 88.02 88.76 93.14 28.24
UIUNet [33] 79.76 78.74 98.69 4.41 65.20 64.57 92.33 13.45 88.86 88.11 98.40 23.17
MTUNet [51] 78.41 79.09 98.02 8.66 67.04 68.24 93.18 12.86 87.39 87.34 98.15 29.43
SCTransNet [52] 80.88 79.45 96.48 6.48 71.22 70.11 94.45 13.94 92.39 92.32 97.71 28.76
SeRankDet [55] 81.26 79.66 98.74 2.77 74.45 70.03 95.12 10.65 94.16 93.34 98.71 22.36
$\star$ DCGANet (Ours) 82.03 79.81 98.88 2.99 75.59 72.16 96.64 9.21 94.98 94.46 98.75 20.34
IV-B Comparison With State-of-the-Arts
Our DCGANet method is rigorously compared against state-of-the-arts methods across three datasets using metrics include IoU, nIoU, Pd, and Fa. The BS-based methods NWMTH [12], FKRW [47]. The HVS-based Methods MPCM [17], WLDM [48], RLCM [49], ILCM [50], TLLCM [41], GSWLCM [18]. The LRM-based methods IPI [42], RIPT [24], PSTNN [26]. The deep learning methods ACM [32], AGPCNet [38], DNANet [34], ISNet [31], UIUNet [33], MTUNet [51], SCTransNet [52], SeRankDet [55]. The experimental results summarized in Tab. III, including three critical observations:
1) Our proposed DCGANet exhibits superior performance compared to state-of-the-art methods across multiple datasets. Specifically designed to tackle the challenges posed by low signal-to-noise ratio (SNR) and blurred target boundaries in infrared small target detection, our approach incorporates a multi-modal feature extractor that precisely extracts target features while effectively suppressing background clutter. Additionally, a dedicated feature fusion module is included to aggregate target details from various hierarchical levels. This innovative design substantially enhances detection accuracy and robustness, resulting in significant advancements in the field of infrared small target detection.
2) The experimental results reveal a performance gap exceeding 15% in Intersection over Union (IoU) between Deep Learning-based methods and traditional approaches, highlighting the inherent limitations of model-driven techniques that rely on global sparsity or contrast priors. In contrast, data-driven DL methods demonstrate superior generalization capabilities by leveraging semantic feature extraction and deeper hierarchical representations. This allows them to significantly reduce sensitivity to hyperparameters in unstructured scenarios, resulting in more robust and adaptable performance.
3) On the NUAA-SIRST dataset, DCGANet surpasses SeRankDet across all evaluation metrics except for the False alarm rate (Fa). While SeRankDet shows moderate effectiveness in combining global and local features, its performance significantly deteriorates in cluttered environments. In contrast, DCGANet addresses these limitations by leveraging precise target localization and dynamic feature fusion, resulting in superior pixel-level segmentation accuracy.
Furthermore, Fig. 6 illustrates receiver operating characteristic (ROC) curves for three datasets. The analysis corroborates the numerical results, confirming DCGANet’s stability and superiority over conventional techniques.
(a) NUAA-SIRST
(b) IRSTD-1K
(c) NUDT-SIRST
Figure 6: ROC curves of different methods on three datasets. Our DCGANet achieves the highest Pd at very low Fa.
IV-C Visualization Analysis
To qualitatively validate the advantages of DCGANet, we analyze representative samples from the three datasets (Fig. 7 - 8). Traditional model-driven methods, which rely on handcrafted features, exhibit high false alarms rates (Fig. 7, yellow boxes) and frequent missed detections (Fig. 8, cyan boxes). In contrast, deep learning-based methods leverage learned hierarchical representations to distinguish targets from complex background clutter effectively, showcasing superior detection robustness.
Gradient-weighted class activation mapping (GradCAM) visualizations for DCGANet (Fig. 9) highlight the focus for each module during the target extraction. The Selective Variable Convolution (SVC) employs diverse convolution forms compared to the baseline UNet, enabling robust feature preservation and edge refinement. The dynamic content guided attention (DCGA) module adaptively refines the distinction between target and background, demonstrating enhanced precision in focusing on true targets compared to static attention mechanisms. Additionally, the Adaptive Dynamic Feature Fusion (ADFF) module improves edge preservation, achieving sub-pixel accuracy in localization. These results collectively demonstrate that DCGANet’s modular, with each component exhibiting clear and discriminative attention patterns.
Figure 7: Visualization comparison of detection results via different methods on representative images from NUAA-SIRST, IRSTD-1K, and NUDT-SIRST datasets, indicate the land forests, oceans skys, and city building interfere. The red, green, and yellow boxes denote correct detection, false alarms, and missed detections, respectively.
Figure 8: Visualization comparesion the detection results of different methods (The red, green, and yellow boxes denote correct detection, false alarms, and missed detections). 3D mesh plot highlighting the features of small targets in the image.
IV-D Ablation Study
IV-D1 Model-wise performance
Ablation experiments were conducted to evaluate the proposed SVC, DCGA, and ADFF modules, demostrating their individual and integrated impacts on model performance, as shown in Tab. IV. The baseline U-Net was used for comparison analysis.
  1. 1.
    Independent efficacy: Each module yields distinct performance improvements, corroborating its theoretical motivations (e.g. SVC’s 5.50% IoU gain on IRSTD-1K and DCGA’s 7.83% improvement).
  2. 2.
    Convolutional Advancement: Replacing vanilla convolutions with the SVC module elevates IoU by 4.62% on NUAA-SIRST and 5.50% on IRSTD-1K, demonstrating its superiority feature encoding multi-scale target features and suppressing background clutter.
  3. 3.
    Attention Mechanism Superiority:The DCGA outperforms established attention modules (e.g. SE, DAM, and CBAM) by adaptively modeling channel-specific spatial importance maps (SIMs) critical for sub-pixel infrared targets.
  4. 4.
    Contextual Fusion:The ADFF module’s dynamic feature fusion strategy, enhances detection accuracy, achieving a 9.62% improvement over baseline fusion strategies on the IRSTD-1K dataset. This demonstrates the effectiveness of leveraging broader contextual information for target detection in complex infrared scenes.
Furthermore, Fig. 9 presents a qualitative comparison of detection results under different model configurations. The visualizations include original infrared inputs, ground-truth masks, attention maps from the baseline U-Net, and results from the progressively integrated DCGANet, showcasing the incremental benefits of each module in target localization and feature enhancement.
Figure 9: Qualitative comparison of detection results on infrared images. From left to right: original images, predictions by U-Net, U-Net+SVC, U-Net+SVC+DCGA, U-Net+SVC+DCGA+ADFF/DCGANet, and ground-truth masks. The improvements in target localization and detail are evident with the addition of each module.
TABLE IV: Quantitative Ablation Study on DCGANet (selective variable convolution is marked SVC, dynamic content guided attention is marked DCGA, adaptive dynamic feature fusion module is marked ADFF), The $\uparrow$ indicates that higher values are better.
Strategy Module NUAA-SIRST IRSTD-1K NUDT-SIRST
SVC DCGA ADFF IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$
  (a) 75.12 75.35 64.42 62.54 89.81 89.58
(b) 79.74 78.53 71.92 69.51 92.42 92.39
(c) 80.11 78.75 72.25 70.21 93.02 92.88
(d) 81.34 79.01 74.86 72.42 93.98 93.52
(e) $\checkmark$ $\checkmark$ $\checkmark$ 82.03 79.81 75.59 72.16 94.98 94.46
IV-D2 Ablation of SVC module
The SVC module integrates deformable convolutio (DC) and multi-dilation convolution (MDC) alongside Standard Convolution to generate more robust feature representations. Ablation results in Tab. V reveal that while DC alone provides a moderate performance gain, the integration of MDC yields significantly greater improvementds. This demonstrates the superior efficacy of the SVC composite convolutional block, where multi-rate dilated convolutions play a pivotal role in expanding receptive fields and capturing hierarchical contextual information.
These findings underscore the critical importance of broad receptive fields in IRSTD, serving as a key motivator for the design of the ADFF module. By leveraging broader spatial context, the ADFF module is engineered to refine target detection further, enhancing both accuracy and robustness in complex infrared scenes. This synergy between the SVC’s multi-convolutional feature extraction and the ADFF’s contextual fusion highlights the importance of hierarchical receptive field design in overcoming the challenges of low SNR and ambiguous target boundaries.
TABLE V: Ablation experiments of convolution branches of SVC.
Strategy Module NUAA-SIRST IRSTD1K NUDT-SIRST
SC MDC DC IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$
(a) 80.35 77.59 72.14 70.12 91.21 91.96
(b) 81.72 78.47 74.01 70.85 93.30 93.88
(c) 80.97 78.22 73.45 70.36 92.64 93.37
(d) 82.03 79.81 75.59 72.16 94.98 94.46
IV-D3 Ablation of DCGA module
Traditional attention modules (SE/DAM/CBAM) fail to address the challenges specific to IRSTD:
  1. 1.
    Under the condition of low SNR, single-channel simulation cannot adequately represent cross-channel variance.
  2. 2.
    Compared with the serial design, the channel shuffling architecture of parallel processing DCGA improves the small target recognition rate.
As shown in Tab. VI, in detection performance, DCGA has significantly improved compared with SE, DAM and CBAM. This shows that DCGA is able to better recalibrate the feature distribution by learning Spatial Importance Maps (SIM) to focus on signal differences between channel-wise, which is particularly critical for enhancing infrared small target features.
TABLE VI: Ablation of DCGA compared with other mainstream attention mechanisms
Strategy NUAA-SIRST IRSTD1K NUDT-SIRST
IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$
SE 78.74 77.90 68.13 68.47 90.93 90.98
DAM 79.88 78.01 70.19 69.11 91.29 91.65
CBAM 80.00 77.88 71.20 68.59 92.30 92.04
DCGA 82.03 79.81 75.59 72.16 94.98 94.46
TABLE VII: Ablation study of DCGA. All the experiments are conducted on three datasets.
Method DCGAcas DCGAw/o r DCGAw/o CS DCGA
parallel CA&SA
cascaded CA&SA
refine ✓(remove CS)
NUAA-SIRST IoU 80.12 75.15 78.54 82.03
nIoU 78.71 75.24 77.12 79.81
IRSTD-1K IoU 72.25 64.41 71.34 75.59
nIoU 70.20 62.97 69.04 72.16
NUDT-SIRST IoU 93.04 89.14 91.19 94.48
nIoU 92.78 89.43 91.01 94.46
Further ablation studies on the DCGA design settings are as follows:
  1. 1.
    DCGAcas represents the cascaded connection of channel attention (CA) and spatial attention (SA) in contrast to the parallel mode of DCGA;
  2. 2.
    DCGAw/o r denotes the parallel connection of channel attention (CA) and spatial attention (SA) but without the refinement process;
  3. 3.
    DCGAw/o CS indicates the parallel connection of channel attention (CA) and spatial attention (SA) but excludes the Channel Shuffle (CS) operation in the refinement stage;
  4. 4.
    DCGA is the proposed method that integrates parallel connection of channel attention (CA) and spatial attention (SA), along with a complete refinement process that includes channel shuffle (CS).
The results are presented in Tab. VII, which reveals that model performance slightly declines when the channel shuffle (CS) operation is removed. This result indicates that CS plays a crucial role in facilitating cross-channel information interaction to highlight small target features. Additionally, the channel attention (CA) and spatial attention (SA) mechanisms in DCGA are deployed in parallel, efficiently modeling the salient regions of small targets by generating initial spatial importance maps (SIMs) in a parallel manner. We also experimented with a cascaded design, and the results show that both parallel and cascaded designs exhibit very similar performance in terms of signal-to-noise ratio (SNR) gain. However, when the refinement process is omitted (DCGAw/o CS), performance significantly deteriorates, further confirming the importance of the refinement stage in enhancing the feature representation of infrared small targets.
V Conclusion
In this paper, we propose a dynamic content guided attention multiscale feature aggregation network (DCGANet), which adopts a “coarse-to-fine” attention principle to achieve high detection accuracy. First, the selective variable convolution (SVC) module is introduced, integrating standard convolution, deformable convolution, and multi-rate dilated convolution. This module expands the receptive field and enhances non-local features, thereby effectively improving the discrimination between targets and backgrounds. Second, the core component of DCGANet is the dynamic content guided attention (DCGA) module, which uses two-stage attention to first direct the network’s focus to significant regions in feature maps and then determine whether these regions correspond to targets or background interference. By retaining the most salient responses, this mechanism suppresses false alarms effectively. Additionally, the adaptive dynamic feature fusion (ADFF) module is proposed to substitute for static feature cascading. The dynamic feature fusion strategy enables DCGANet to adaptively integrate contextual features, enhancing its capability to discriminate between true targets and false alarms. DCGANet establishes new benchmarks across multiple datasets.

### v2

@@
-Selective Variable Convolution Meets Dynamic Content Guided Attention for Infrared Small Target Detection
+Selective Variable Convolution Meets Dynamic Content-Guided Attention for Infrared Small Target Detection
 Abstract
@@
 Due to the convolution operation in Convolutional Neural Networks (CNNs), applying traditional CNNs to IRSTD presents challenges, since the feature extraction of small targets is often insufficient, resulting in the loss of critical features.
-To address these issues, we propose a dynamic content guided attention multiscale feature aggregation network (DCGANet), which adheres to the attention principle of ’coarse-to-fine’ and achieves high detection accuracy.
+To address these issues, we propose a dynamic content-guided attention multiscale feature aggregation network (DCGANet), which adheres to the attention principle of ’coarse-to-fine’ and achieves high detection accuracy.
 First, we propose a selective variable convolution (SVC) module that integrates the benefits of standard convolution, irregular deformable convolution, and multi-rate dilated convolution.
-This module is designed to expand the receptive field and enhance non-local features, thereby effectively improving the discrimination of targets from backgrounds.
-Second, the core component of DCGANet is a two-stage content guided attention module.
-This module employs two-stage attention mechanism to initially direct the network’s focus to salient regions within the feature maps and subsequently determine whether these regions correspond to targets or background interference.
+This module is designed to expand the receptive field and enhance non-local features, thereby effectively improving the discrimination between targets and backgrounds.
+Second, the core component of DCGANet is a two-stage content-guided attention module.
+This module employs a two-stage attention mechanism to initially direct the network’s focus to salient regions within the feature maps and subsequently determine whether these regions correspond to targets or background interference.
 By retaining the most significant responses, this mechanism effectively suppresses false alarms.
-Additionally, we propose adaptive dynamic feature fusion (ADFF) module to substitute for static feature cascading.
+Additionally, we propose an Adaptive Dynamic Feature Fusion (ADFF) module to substitute for static feature cascading.
 This dynamic feature fusion strategy enables DCGANet to adaptively integrate contextual features, thereby enhancing its ability to discriminate true targets from false alarms.
@@
 Index Terms:
-Infrared small target, selective variable convolution, dynamic content guided attention, adaptive dynamic feature fusion.
+Infrared small target , Selective variable convolution ,Dynamic content-guided attention ,Adaptive dynamic feature fusion
 I Introduction
 Infrared imaging technology, a crucial component of infrared monitoring systems, excels at capturing thermal signatures of targets and demonstrates robust performance in low-light and foggy environments.
-In military [1], ocean rescue [2], and environmental monitoring [3] applications, targets of interest at long imaging distances often appear as small spots [3].
+In military [1], ocean rescue [2], and environmental monitoring [3, 4] applications, targets of interest at long imaging distances often appear as small spots [3].
 Consequently, infrared small target detection (IRSTD) is essential for extracting valuable information from complex backgrounds.
-Figure 1: The complex background faced by image target detection, the red and green bounding boxes are zoom in area: (a)Dense target of different types such as pedestrians and cars, (b)Target occlusion by other objects such as trees, (c)Strong lighting conditions of sunlight, and (d) Dim lighting conditions during nights.
+Figure 1: Due to the long imaging distance, real targets (red boxes) lack clear texture and structure, appearing as tiny bright spots.
+Meanwhile, high-brightness clutter in the background, such as from clouds, buildings, or the ground (green boxes), actively mimics real targets in local morphology, size, and brightness, creating indistinguishable interference.
 Despite the critical role of IRSTD in infrared warning systems, as illustrated in Fig.
@@
 Addressing these challenges is essential for achieving robust and precise detection in infrared warning systems.
-It is imperative to suppress background clutter by enhancing the extraction of inherent target features and integrating contextual analysis, which is pivotal to improving IRSTD performance.
-Traditional IRSTD methodologies can be categorized into three primary approaches.
+To improve IRSTD performance, it is crucial to suppress background clutter.
+This can be achieved by enhancing the extraction of inherent target features and integrating contextual analysis.
+Traditional data-driven IRSTD methodologies can be categorized into three primary approaches.
 They are Background Suppression-based (BS-based), Human Visual System-based (HVS-based), and Low Rank Matrix-based (LRM-based) methods.
-The BS-based methods [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] utilize specialized filters to extract small targets.
-The HVS-based methods [16, 17, 18, 19, 20, 21, 22, 23] identify small targets by exploiting local information features.
-The LRM-based methods [24, 25, 26, 27] detect small targets through the decomposition and subsequent reconstruction of infrared images into sparse matrices.
+The BS-based methods [6, 7, 8, 9] utilize specialized filters to extract small targets.
+The HVS-based methods [10, 11, 12] identify small targets by exploiting local information features.
+The LRM-based methods [13, 14] detect small targets through the decomposition and subsequent reconstruction of infrared images into sparse matrices.
 However, these traditional methods are limited in their ability to learn from datasets or dynamically adjust parameters in response to changing environmental conditions, leading to suboptimal performance in complex scenarios.
-Recent advances in deep learning have yielded substantial improvements for IRSTD.
-Previous research in deep learning methods emphasized weak feature expressions, such as gradient edge [28, 29, 30, 31], contextual information [32, 33, 34, 35, 36], and local contrast [37, 38, 39].
+Recent advances in deep learning have yielded substantial improvements for data-driven IRSTD [15].
+Previous research in deep learning methods emphasized weak feature expressions, such as gradient edge [16], contextual information [17, 18, 19, 20], and local contrast [21, 22].
 Despite their progress, fundamental challenges in IRSTD remain unaddressed.
@@
     Static Attention for Interference Suppression: Traditional convolution and pooling strategies often struggle to effectively separate background and target features, while the saliency of infrared targets may be distributed in specific channels (such as thermal radiation intensity, gradient changes, etc.).
-However, existing attention mechanisms (such as SE [10], CBAM [34]) model channel relationships statically or at a single scale, causing target information to be masked by background signals, reducing detection sensitivity.
+However, existing attention mechanisms (such as SE [23], CBAM [19]) model channel relationships statically or at a single scale [20], causing target information to be masked by background signals, reducing detection sensitivity.
   3. 3.
@@
 These limitations collectively impede robust detection in real-world scenarios, where targets exhibit dynamic scales and unpredictable noise patterns.
-Facing the challenge of IRSTD, existing convolutional neural networks often struggle to effectively address the feature extraction of infrared small targets.
-To overcome these challenges, there is an imperative need for a novel framework, which is derived from three pivotal observations: (1) Infrared small targets require hierarchical dynamic receptive fields to jointly capture edge sharpness and thermal radiation context; (2) Feature fusion should be selectively guided by high-level semantics to suppress noise propagation; (3) Attention mechanisms must incorporate infrared physical priors and consider input channels according to wavelength sensitivity, while adapting to environmental dynamics.
-Therefore, our framework should not only effectively extract subtle features of infrared small targets through advanced feature extraction techniques but also dynamically optimize multi-stage feature fusion processes.
-This framework requires a deep understanding of the background environment, rather than simply relying on local image information near the target.
-To address the aforementioned challenges, we propose DGCANet, which consists of three core innovations.
+Facing the challenges encountered in existing infrared small target detection, this study sets its core objective: to construct a high-performance detection framework capable of synergistically optimizing feature learning, contextual awareness, and feature fusion.
+Firstly, a dynamic feature extraction mechanism is introduced.
+This mechanism aims to resolve the inherent conflict between local receptive fields and global contextual modeling, adaptively adjusting its information aggregation scope to enhance the representational capacity of foundational features.
+Secondly, a multi-scale context-guided attention strategy is employed.
+This strategy first expands the perceptual range of local features and conducts intelligent grouping and refinement along the channel dimension based on the physical properties of infrared imaging, thereby generating multi-scale attention maps.
+Finally, to address missed detections and false alarms caused by improper fusion, an adaptive feature fusion method is proposed.
+This method dynamically evaluates and balances the contributions of features from different hierarchical levels, intelligently integrating complementary information.
+In summary, this research is committed to achieving breakthroughs in three critical aspects—feature extraction, information perception, and feature fusion—through the above synergistic approaches.
+The ultimate goal is to establish a robust infrared small target detection system capable of delivering high-precision and high robustness in complex, dynamic infrared scenarios.
+To address the aforementioned challenges, we propose DCGANet, which consists of three core innovations.
 First, to improve the recognition of real targets, suppress background noise, and discriminate between them and false alarms, we proposed selective variable convolution (SVC), which consists of multi-branch adaptive kernels.
-The SVC resolving the local-global trade-off, which can accurately extract features of weak targets and significantly reduce background noise.
-Second, a dynamic content guided attention(DCGA) mechanism was proposed.
-The DGCA first performs local feature extraction and expanded receptive field extraction using parallel ordinary convolution and dilated convolution.
-To generate multi-scale coarse-grained attention maps, through channel grouping refinement strategy, the input feature channels are grouped according to physical meanings (such as thermal radiation intensity and gradient texture), and captures insensitive features of weak targets through fine attention allocation.
-Last, we proposed Adaptive Dynamic Feature Fusion (ADFF) to simulate the attention regulation mechanism of the human vision systems(HVS).
-The integration of ADFF with a wide receptive field significantly improves the distinguish between real targets and false positives.
+The SVC resolves the local-global trade-off, which can accurately extract features of weak targets and significantly reduce background noise.
+Second, we proposed a Dynamic Content-Guided Attention (DCGA) mechanism.
+This mechanism simulates the "coarse-to-fine" search pattern of the human visual system.
+Through a two-stage process, it dynamically generates a dedicated Spatial Importance Map (SIM) for each channel of the feature map, thereby precisely enhancing target features while suppressing background interference.
+Last, to address the issue of improper fusion between different hierarchical features, we designed the Adaptive Dynamic Feature Fusion (ADFF) module.
+It replaces traditional skip connections, intelligently weighing and merging features from the encoder and decoder through a dynamic spatial selection mechanism.
+This effectively suppresses the contamination of deep semantic features by shallow-level noise.
 The main contributions of this paper are listed as follows:
   1. 1.
-    We propose DCGANet, a novel framework that unifies feature learning, dynamic content guided attention, and feature fusion in a hierarchical dynamic pipeline.
+    We propose DCGANet, a novel framework that unifies feature learning, contextual guide attention, and feature fusion in a hierarchical dynamic pipeline.
 This framework bridges the gap between rigid CNN designs and the dynamic, sparse infrared target, achieving joint optimization of accuracy, robustness, and interpretability.
@@
 Model-driven methods formalize IRSTD as anomaly recognition against backgrounds, leveraging three computational paradigms: filters-based, human visual system (HVS) simulations-based, and sparse matrix reconstruction.
-The filter-based method [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] mainly focuses on suppressing background clutter and highlighting targets while suppressing isolated noise through filter stacking.
+The filter-based method [6, 7, 8] mainly focuses on suppressing background clutter and highlighting targets while suppressing isolated noise through filter stacking.
 For instance, multi-scale Top-Hat [9, 25] filtering algorithm, which combines multi-scale transformation, can improve detection performance.
 However, when the infrared image contains complex backgrounds, the algorithm will exhibit a lower signal-to-noise ratio (SNR).
-To further improve the performance, the Human Visual System-Based (HVS-Based) methods [16, 17, 18, 19, 20, 21, 22, 23, 41] utilize the principle of human visual contrast to suppress background brightness and enhance local information of infrared small targets, such as LCM [16], and TLLCM [41].
+To further improve the performance, the Human Visual System-Based (HVS-Based) methods [10, 11, 12, 26] utilize the principle of human visual contrast to suppress background brightness and enhance local information of infrared small targets, such as LCM [10], and TLLCM [27].
 However, HVS-based methods still struggle to maintain detection stability when faced with high interference from bright edges or points.
-From another perspective, sparse matrix reconstruction treats small targets as sparse matrices and backgrounds as low-rank matrices [24, 25, 26, 27, 42], detecting small targets by decomposition low-rank matrices.
+From another perspective, sparse matrix reconstruction treats small targets as sparse matrices and backgrounds as low-rank matrices [13, 14, 28], detecting small targets by decomposition low-rank matrices.
 For instance, the Infrared Patch Image (IPI) [29] model divides infrared images into blocks and recombines them into low-rank sparse matrices.
@@
 Although deep learning methods have made substantial progress, they still face challenges such as over-parameterized architectures and inflexible receptive fields [30], impending comprehensive capture of the details of small targets.
-Excessive reliance on local information [31] also limits their ability to model global contextual features, leading to high false alarm rates in noisy scenarios.
+Excessive reliance on local information [32, 33] also limits their ability to model global contextual features, leading to high false alarm rates in noisy scenarios.
 DCGANet distinguishes itself through two core innovations:
@@
   2. 2.
-    HVS-mimetic dynamic attention: The DCGA module implements ’coase-to-fine’ saliency mapping, mirroring human visual system search patterns through channel-specific attention refinement.
+    HVS-mimetic dynamic attention: The DCGA module implements ’coarse-to-fine’ saliency mapping, mirroring human visual system search patterns through channel-specific attention refinement.
 II-B Vision Attention Mechanisms on IRSTD
@@
 Despite these advancements, existing methods still lack robust global feature understanding in IRSTD.
-In contrast, dynamic content guided attention (DCGA) implements a coarse-to-fine attention cascade: it first generates a rough spatial attention map, then refines it into a fine-grained map by integrating spatial and channel attention.
+In contrast, Dynamic Content-Guided Attention (DCGA) implements a coarse-to-fine attention cascade: it first generates a rough spatial attention map, then refines it into a fine-grained map by integrating spatial and channel attention.
 This process refines initial spatial attention maps (SIM) for each channel to focus on the most representative features, enhancing the recognition of suspected small targets.
@@
 This dual optimization not only boosts small target detection accuracy but also strengthens false alarm suppression capabilities.
-III Method
-III-A Overall Architecture
-Figure 2: Overall architecture of the proposed DCGANet, built upon a U-Net-like backbone and integrating the key Selective Variable Convolution (SVC), Dynamic Content Guided Attention (DCGA), and Adaptive Dynamic Feature Fusion (ADFF) modules.
-The DCGANet architecture, depicted in Fig.
-2, is constructed upon a U-Net-like framework and comprises three components: selective variable convolution (SVC), dynamic content guided attention (DCGA), and adaptive dynamic feature fusion (ADFF) modules.
-Specifically, the encoder path consists of three cascaded layers integrating SVC and DCGA modules in sequence.
-The SVC replaces standard convolutions to enhance subtle target signatures while expanding receptive fields for augmented context modeling.
-The DCGA mechanism generates channel-specific spatial importance maps (SIMs) through input-guided attention calibration, enabling per-channel saliency emphasis critical for sub-pixel infrared targets.
-On the decoder path, two ADFF modules dynamically merge multi-scale encoder features, replacing the original skip-connections and achieving context-aware noise suppression through three-stage fusion.
-III-B Selective Variable Convolution
-The limited receptive field of traditional convolutional layers in CNNs hinders their ability in IRSTD.
-To address this, we propose the selective variable convolution (SVC) module, a robust basic block aimed at significantly enhancing the network’s feature learning ability.
-As shown in Fig.
-3, the SVC combines synergy of the tri-branch features.
-The first branch employs standard convolution(SConv) to preserve local spatial feature extraction, retaining elementary patterns of small targets and maintaining the stability of fundamental detection performance.
-The second branch utilizes deformable convolution (DConv), which has been proven effective in edge delineation for irregularly shaped targets, enhancing target features across diverse scales.
-The third branch, multi-dilation convolution (MDConv), utilizes multi-rate dilated kernels (rates = 2, 4, 8) to extend the receptive field, thereby aggregating multi-scale contexts to suppress structured clutter.
-The SVC enables accurate extraction of weak target features and significant reduction of background noise.
+III Methodology
+Figure 2: The overall architecture of the proposed DCGANet.
+It is built upon a U-Net-like backbone and integrates the key Selective Variable Convolution (SVC), Dynamic Context-Guided Attention (DCGA), and Adaptive Dynamic Feature Fusion (ADFF) modules.
+Figure 3: The structure of the SVC module: a parallel branch architecture that performs different types of convolution operations on the input feature map.
+This design enables the encoding stage to extract features at different scales and preserve details.
+The overall architecture of DCGANet is illustrated in Figure 2, built upon a U-Net-like framework and composed of three core components: the Selective Variable Convolution (SVC), Dynamic Context-Guided Attention (DCGA), and Adaptive Dynamic Feature Fusion (ADFF) modules.
+Specifically, the encoder path consists of three cascaded layers, each sequentially integrating an SVC and a DCGA module.
+SVC replaces standard convolution to enhance faint target features while expanding the receptive field for stronger contextual modeling.
+The subsequent DCGA mechanism generates channel-specific Spatial Importance Maps (SIMs) through input-guided attention calibration, enabling per-channel saliency emphasis crucial for sub-pixel infrared targets.
+On the decoder path, two ADFF modules dynamically fuse multi-scale features from the encoder, replacing the original skip connections and effectively suppressing noise through a context-aware fusion strategy.
+III-A Selective Variable Convolution
+In Convolutional Neural Networks (CNNs), the limited receptive field of traditional convolutional layers hinders their ability to capture global contextual information in Infrared Small Target Detection (IRSTD).
+To address this, we propose the Selective Variable Convolution (SVC) module, a powerful foundational block designed to significantly enhance the network’s feature learning capabilities.
+The structure of SVC, shown in Figure 3, enhances expressive power by integrating features from three parallel branches.
+  • Standard Convolution (SConv): The first branch employs a standard convolution to preserve local spatial features, maintaining the stability of basic patterns for small targets and foundational detection performance.
+  • Deformable Convolution (DConv): The second branch utilizes deformable convolution, which excels at outlining the edges of irregularly shaped targets, enhancing target features at different scales.
+  • Multiple Dilation Convolution (MDConv): The third branch is a multiple dilation convolution, which uses kernels with different dilation rates (rates = 2, 4, 8) to expand the receptive field, thereby aggregating multi-scale context to suppress structured clutter.
+The mathematical expressions for these three branches are as follows:
 $X_{s}$ $=\operatorname{SConv}(X_{\text{in}}),$ (1)
 $X_{c}$ $=\operatorname{DConv}(X_{\text{in}}),$ (2)
-$X_{m}$ $=\operatorname{MDConv}(X_{\text{in}}).$ (3)
-where $X_{\text{in}}$ is the input feature, $X_{s}$, $X_{c}$, and $X_{m}$ are the outputs of SConv, DConv, and MDConv, respectively.
-The output of these branches is fused through channel-wise concatenation and a $1\times 1$ convolution, yielding consolidated representations enriched in both local precision and global discriminability.
+$X_{m}$ $=\operatorname{MDConv}(X_{\text{in}}),$ (3)
+where $X_{\text{in}}$ is the input feature, and $X_{s}$, $X_{c}$, and $X_{m}$ are the outputs of the three branches, respectively.
+The outputs of these branches are fused via channel-wise concatenation and a $1\times 1$ convolution, yielding an integrated representation enriched in both local precision and global differentiability.
 $X_{\text{out}}=\operatorname{Conv}_{1\times 1}\left(\operatorname{Concat}\left(X_{s},X_{c},X_{m}\right)\right),$ (4)
-where $X_{\text{out}}$ is the output of SVC module, $X_{s}$, $X_{c}$, and $X_{m}$ are the outputs of SConv, DConv, and MDConv, respectively.
-By enhancing the extraction of edges and details through a multi-convolution strategy, SVC effectively boosts recognition capability in IRSTD.
-The expanded receptive field and deformable feature extraction offer substantial advantages in contextual perception and noise suppression, thereby improving the clarity and reliability of feature extraction.
-Figure 3: The structure of the SVC: a parallel branching architecture to perform different types of convolution operations on the input feature maps.
-This design enables the encoding phase to extract features and preserve details at various scales.
-III-C Dynamic Content Guided Attention
-While the SVC module expands the receptive field and enhances non-local feature extraction, it still struggles to identify the target when dealing with background clutter.
-Previous attention mechanisms, such as the static channel-wise or spatial weight separation mechanisms of SE [10], CSAM [34], face two key issues.
-First, there is a lack of information interaction between channel-wise and spatial attention.
-Second, single-channel spatial weights cannot adapt to the diverse distribution of multiple feature channels and pose a risk of information loss due to the dilution of key target features.
-To address these challenges, we propose the dynamic content guided attention(DCGA) module, which uses a coarse-to-fine approach to generate self-attention maps(SIM) for specific channels, effectively simulating the ability of HVS to focus on highlighting features and suppress false alarms.
-The DCGA is shown in Fig.
-4.
-The first stage includes coarse-grained attention with interaction between channel-wise and spatial attention, while the second stage includes fine-grained calibration with dynamic SIM generation, obtaining a unique SIM for each single-channel input features in a coarse-to-fine manner.
-We first calculate the corresponding Wc and Ws by following.
-III-C1 Stage I: Coarse-Grained Attention with Joint Channel-Space Fusion
-The stage include Channel attention, Spatial Attention, and joint attention.
-Channel attention: A dual-bottleneck structure extracts global channel responses by aggregating spatial information through global average pooling:
-$W_{c}=\operatorname{\mathcal{C}}_{1\times 1}\left(\operatorname{ReLU}\left(\operatorname{\mathcal{C}}_{1\times 1}\left(X_{GAP}^{\text{c}}\right)\right)\right)$ (5)
-where $X_{GAP}^{\text{c}}$ denotes channel-wise squeezed features from global average pooling.
-Spatial Attention: Spatial saliency is captured by concatenating multi-perspective features from global average pooling (GAP) and global max pooling (GMP):
-$W_{s}=\operatorname{\mathcal{C}}_{7\times 7}\left(\operatorname{Concat}\left(X_{GAP}^{\text{s}},X_{GMP}^{\text{s}}\right)\right).$ (6)
-where $X_{GAP}^{\text{s}},X_{GMP}^{\text{s}}$ represent spatial descriptors from global average pooling (GAP) and global max pooling (GMP).
-Interaction Attention: we fuse $W_{c}$ and $W_{s}$ together using an element-wise addition operation following broadcast rules to obtain a rough SIM $W_{\text{coa}}$.
+where $X_{\text{out}}$ is the final output of the SVC module.
+Through this multi-convolution strategy, SVC effectively improves the recognition of target edges and details, and its expanded receptive field and deformable feature extraction provide significant advantages for context-aware modeling and noise suppression.
+III-B Dynamic Context-Guided Attention
+Although the SVC module enhances feature extraction, it may still face difficulties when dealing with complex background clutter.
+Traditional attention mechanisms suffer from two key problems: first, a lack of information interaction between channel and spatial attention; and second, a single spatial weight for all channels cannot adapt to the diverse distributions across multiple feature channels, risking the dilution of key target information.
+To address these challenges, we propose the Dynamic Context-Guided Attention (DCGA) module.
+The structure of DCGA, shown in Figure 4, employs a coarse-to-fine approach to Spatial Importance Map (SIM), effectively simulating the Human Visual System’s (HVS) ability to focus on salient features and suppress false positives.
+Figure 4: The structure of the DCGA module: It employs a parallel approach to equally weight spatial and channel information.
+It also incorporates a channel shuffling of channel and spatial information to enhance target segmentation capabilities.
+III-B1 Stage 1: Coarse-Grained Attention and Joint Channel-Spatial Fusion
+This stage generates a preliminary attention map.
+Channel Attention: A dual-bottleneck structure is used, aggregating spatial information via Global Average Pooling (GAP) to extract the global channel response $W_{c}$:
+$W_{c}=\operatorname{\mathcal{C}}_{1\times 1}\left(\operatorname{ReLU}\left(\operatorname{\mathcal{C}}_{1\times 1}\left(X_{GAP}^{\text{c}}\right)\right)\right),$ (5)
+where $X_{GAP}^{\text{c}}$ is the channel-compressed feature obtained from Global Average Pooling.
+Spatial Attention: Spatial saliency $W_{s}$ is captured by concatenating multi-view features from Global Average Pooling (GAP) and Global Max Pooling (GMP):
+$W_{s}=\operatorname{\mathcal{C}}_{7\times 7}\left(\operatorname{Concat}\left(X_{GAP}^{\text{s}},X_{GMP}^{\text{s}}\right)\right),$ (6)
+where $X_{GAP}^{\text{s}}$ and $X_{GMP}^{\text{s}}$ represent the spatial descriptors from GAP and GMP, respectively.
+Interactive Attention: Following broadcasting rules, we fuse $W_{c}$ and $W_{s}$ using element-wise addition to obtain a coarse SIM $W_{\text{coa}}$.
 $W_{\text{coa}}=W_{c}\oplus W_{s},$ (7)
-where $W_{\text{coa}}\in\mathbb{R}^{C\times H\times W}$ is rough SIM, $\oplus$ denotes element-wise addition with automatic dimension expansion.
-Figure 4: The proposed structure of the DCGA: employs a parallel approach to equally compute the weights of spatial and channel information.
-It also incorporates random mixing of channel and spatial information to enhance the capability for target segmentation.
-III-C2 Stage II: Fine-Grained Calibration with Dynamic SIM Generation
-The fine-grained calibration include channel shuffling fusion and dynamic weight generation.
-Channel Shuffling Fusion: The raw feature $X\in\mathbb{R}^{C\times H\times W}$ and coarse attention map $W_{coa}$ are concatenated and then shuffled along the channel dimension [46], establishing inter-channel correlations while preserving spatial details.
-Dynamic Weight Generation: A gated convolution layer $\mathcal{GC}_{7\times 7}$ followed by sigmoid activation generates channel-specific SIM :
-$W=\sigma\left(\operatorname{\mathcal{GC}}_{7\times 7}\left(\operatorname{CS}\left(\operatorname{Concat}\left(X,W_{\text{coa}}\right)\right)\right)\right)$ (8)
-where $\operatorname{CS}(\cdot)$ denotes the channel shuffling operation.
-This enables content-adaptive calibration by allowing each channel’s SIM to dynamically adjust based on local feature patterns.
-DCGA assigns a unique SIM to each channel, guiding the model to focus on the important areas of each channel.
-Therefore, more useful information encoded in the features can be emphasized to effectively improve IRSTD.
-III-D Adaptive Dynamic Feature Fusion Module
-The proposed Adaptive Dynamic Feature Fusion(ADFF) module aims to enhance the network’s ability to distinguish between actual targets and similar false alarms.
-Specifically, we introduced a dynamic space selection mechanism within the original skip connections, as shown in Fig.
-5.
-This mechanism leverages multi-scale features to enable the network to focus on the areas with the richest spatial contextual information, thereby increasing sensitivity to the target area.
-The ADFF module is specifically designed to process the feature maps of the encoder-decoder pair in stage i, where the feature maps are represented as:
-$X_{i},X^{\prime}_{i}\in\mathbb{R}^{C_{i}\times\frac{H}{2^{i}}\times\frac{W}{2^{i}}}$ (9)
-where $C_{i}$ denotes channel number, $\frac{H}{2^{i}}$ and $\frac{W}{2^{i}}$ are height and width of feature maps.
-Then we applied channel-wise average pooling and max pooling operations to feature maps $X_{i}$and $X^{\prime}_{i}$, represented as $P_{avg}(\cdot)$and $P_{max}(\cdot)$, respectively.
+where $W_{\text{coa}}\in\mathbb{R}^{C\times H\times W}$, and $\oplus$ denotes element-wise addition with automatic dimension expansion.
+III-B2 Stage 2: Fine-Grained Calibration for Dynamic SIM Generation
+This stage refines the coarse attention map.
+Channel-Shuffle Fusion: The original features $X\in\mathbb{R}^{C\times H\times W}$ and the coarse attention map $W_{coa}$ are concatenated and then shuffled along the channel dimension[35], thereby establishing cross-channel correlations while preserving spatial details.
+Dynamic Weight Generation: A gated convolutional layer $\mathcal{GC}_{7\times 7}$ followed by a Sigmoid activation function generates the final channel-specific SIM $W$:
+$W=\sigma\left(\operatorname{\mathcal{GC}}_{7\times 7}\left(\operatorname{CS}\left(\operatorname{Concat}\left(X,W_{\text{coa}}\right)\right)\right)\right),$ (8)
+where $\operatorname{CS}(\cdot)$ denotes the channel shuffle operation and $\sigma(\cdot)$ is the Sigmoid function.
+This content-adaptive calibration allows the SIM for each channel to be dynamically adjusted based on local feature patterns, guiding the model to focus on important regions within each channel.
+Figure 5: The structure of the ADFF module: It integrates the encoding and decoding stages of the network via a spatially adaptive selection mechanism.
+It utilizes cross-channel max-pooling and average-pooling for precise target extraction.
+III-C Adaptive Dynamic Feature Fusion
+Our proposed Adaptive Dynamic Feature Fusion (ADFF) module is designed to enhance the network’s ability to distinguish real targets from similar false positives.
+As shown in Figure 5, we introduce a dynamic spatial selection mechanism into the original skip connection, leveraging multi-scale features to enable the network to focus on regions richest in spatial context.
+The ADFF module is specifically designed to process the encoder features $X_{i}$ and decoder features $X^{\prime}_{i}$ at stage $i$:
+$X_{i},X^{\prime}_{i}\in\mathbb{R}^{C_{i}\times\frac{H}{2^{i}}\times\frac{W}{2^{i}}},$ (9)
+where $C_{i}$ is the number of channels, and $\frac{H}{2^{i}}$ and $\frac{W}{2^{i}}$ are the height and width of the feature maps.
+We first apply channel-wise average-pooling and max-pooling operations to the concatenated feature map:
 $AD_{i}^{\text{avg}}$ $=\operatorname{AvgPool}\left(\operatorname{Concat}\left[X_{i},X^{\prime}_{i}\right]\right),$ (10)
-$AD_{i}^{\text{max}}$ $=\operatorname{MaxPool}\left(\operatorname{Concat}\left[X_{i},X^{\prime}_{i}\right]\right).$ (11)
-where $AvgPool$ is average pooling operation, $MaxPool$ is max pooling operation.
-These two different pooling operations extract the average and maximum feature descriptors of the space, and the generated spatial pooling features are aggregated into a matrix to facilitate information exchange between different spatial descriptors.
-Next, we use the convolutional layer $F_{2\to N}(\cdot)$to transform these aggregated features into $N$spatial attention maps.
-These spatial attention maps can effectively guide the subsequent feature fusion process, highlight salient regions, and suppress redundant background information.
-Figure 5: The proposed structure of ADFF: integrates the encoding and decoding stages of the network through a spatially adaptive selection mechanism.
-It utilizes both maximum and average pooling across channels to achieve precise target extraction.
+$AD_{i}^{\text{max}}$ $=\operatorname{MaxPool}\left(\operatorname{Concat}\left[X_{i},X^{\prime}_{i}\right]\right),$ (11)
+where $\operatorname{AvgPool}$ and $\operatorname{MaxPool}$ are the average-pooling and max-pooling operations, respectively.
+Next, we concatenate these two pooled features and pass them through a convolutional layer and a Sigmoid function to generate the final spatial attention map $AD_{\text{fuse}}$:
 $AD_{\text{fuse}}=\sigma\left(\operatorname{Conv}_{k\times k}\left(\operatorname{Concat}\left(AD_{i}^{\text{avg}},AD_{i}^{\text{max}}\right)\right)\right),$ (12)
-where $\sigma$ denotes the sigmoid function, $\operatorname{Concat}(\cdot)$ is the channel-wise concatenation, and $\operatorname{Conv}_{k\times k}(\cdot)$ represents the convolutional layer (e.g., with a $7\times 7$ kernel, typically outputting a single channel) used to generate the attention map.
-The final output $Y_{i}$ of the ADFF module is obtained by first applying the generated attention map $AD_{\text{fuse}}$ to weight the input features $X_{i}$ and $X^{\prime}_{i}$ via element-wise multiplication, and then fusing the results through element-wise addition.
-This process is formulated as:
-$Y_{i}=X_{i}\odot AD_{\text{fuse}}+X^{\prime}_{i}\odot AD_{\text{fuse}}$ (13)
-where $\odot$ denotes element-wise multiplication, $Y_{i}$ is the output of ADFF.
-The fusion process generates dynamic fusion weights, which can adaptively adjust the importance of each feature map in the final decision.
-Through this adaptive selective fusion strategy, the ADFF module can automatically identify and integrate the most discriminative features, thereby improving the network’s representation ability of IRSTD.
-IV Experiment
+where $\operatorname{Conv}_{k\times k}$ is a convolutional layer with a kernel size of $k\times k$ (e.g., $7\times 7$).
+The final output $Y_{i}$ of the ADFF module is obtained by element-wise multiplication of the generated attention map $AD_{\text{fuse}}$ with the input features $X_{i}$ and $X^{\prime}_{i}$ respectively, and then adding the results:
+$Y_{i}=(X_{i}\odot AD_{\text{fuse}})+(X^{\prime}_{i}\odot AD_{\text{fuse}}),$ (13)
+where $\odot$ denotes element-wise multiplication.
+Through this adaptive selective fusion strategy, the ADFF module can automatically identify and integrate the most discriminative features, thereby enhancing the network’s representation capability in infrared small target detection.
+TABLE I: Configurations for all comparative experiments
 […]
-table
-TABLE I: Configurations for all comparative experiments
-Methods(Source & Year) Key parameters configurations
-Background Suppression Methods
-NWMTH [12]PR’2010 $R_{o}=9,R_{i}=4$
-FKRW [47]TGRS’2019 Windows size: 11, $K=4,p=6,\beta=200$
-HVS-based Methods
-MPCM [17]PR’2016 $N=1,3,\cdots,9$, threshold $k=3$
-WLDM [48]TGRS’2016 $L=4,m=2,n=2$, threshold $k=2$
-RLCM [49]GRSL’2018 $k_{1}=[2,5,9],k_{2}=[4,9,16]$, scale: 3, threshold $k=1$
-ILCM [50]GRSL’2020 Cell size: $3\times 3$, threshold $k=3$
-TLLCM [41]GRSL’2020 Scales:[5,7,9] $\lambda=0.5$, threshold $k=3$
-GSWLCM [18]GRSL’2022 Local Window Structures: $[3,5,7,9]$, $\delta=0.01,k=20$
-Low-rank and Sparse Decomposition
-IPI [42]TIP’2013 Patch size: 50, sliding step: 10, $\lambda=1/\sqrt{\mathrm{min}(m,n)}$, $\epsilon=10^{-7}$
-RIPT [24]JSTARS’2017 Patch size: 50, sliding step: 10, $\lambda=2/\sqrt{\mathrm{min}(m,n)}$, $\epsilon=10^{-2}$, $\omega=10^{-7}$
-PSTNN [26]RS’2019 Patch size: 40, sliding step: 40, $\lambda=0.6/\sqrt{\mathrm{max}(n_{1},n_{2})\times n_{3}}$, $\epsilon=10^{-7}$
-Deep Learning Methods
-ACM [32]WACV’2021 Backbone: FPN, layer blocks: [4, 4, 4], channels: [8, 16, 32, 64]
-AGPCNet [38]TAES’2023 Backbone: resnet34, scales: [10, 6, 5, 3], reduce ratios: [16, 4], gca type: patch, gca att: post
-DNANet [34]TIP’2022 Backbone: resnet18, layer blocks: [2, 2, 2, 2], filter: [16, 32, 64, 128, 256]
-ISNet [31]CVPR’2022 Backbone: resnet18, layer blocks: [4, 4, 4], channels: [8, 16, 32, 64], fuse mode: AsymBi
-UIUNet [33]TIP’2023 Channels: [64, 128, 256, 512], fuse mode: AsymBi
-MTUNet [51]TGRS’2023 Backbone: resnet18, layer blocks: [2, 2, 2, 2], filter: [16, 32, 64, 128, 256]
-SeRankDet TGRS’2024 Channels:[64, 128, 256, 512, 1024],o=3
-SCTransNet [52] TGRS’2024 Backbone:Swin-Tiny, Layer Num=[2,6,6,2], Head Num=[3,6,12,24]
+[…]
+[…]
+FKRW [37] (TGRS’2019) Windows size: 11, $K=4,p=6,\beta=200$
+[…]
+[…]
+WLDM [38](TGRS’2016) $L=4,m=2,n=2$, threshold $k=2$
+RLCM [39](GRSL’2018) $k_{1}=[2,5,9],k_{2}=[4,9,16]$, scale: 3, threshold $k=1$
+ILCM [40](GRSL’2020) Cell size: $3\times 3$, threshold $k=3$
+TLLCM [27](GRSL’2020) Scales:[5,7,9] $\lambda=0.5$, threshold $k=3$
+GSWLCM [12](GRSL’2022) Local Window Structures: $[3,5,7,9]$, $\delta=0.01,k=20$
+[…]
+[…]
+RIPT [13] (JSTARS’2017) Patch size: 50, sliding step: 10, $\lambda=2/\sqrt{\mathrm{min}(m,n)}$, $\epsilon=10^{-2}$, $\omega=10^{-7}$
+PSTNN [28](RS’2019) Patch size: 40, sliding step: 40, $\lambda=0.6/\sqrt{\mathrm{max}(n_{1},n_{2})\times n_{3}}$, $\epsilon=10^{-7}$
+[…]
+[…]
+AGPCNet [22](TAES’2023) Backbone: resnet34, scales: [10, 6, 5, 3], reduce ratios: [16, 4], gca type: patch, gca att: post
+DNANet [19](TIP’2022) Backbone: resnet18, layer blocks: [2, 2, 2, 2], filter: [16, 32, 64, 128, 256]
+ISNet [32](CVPR’2022) Backbone: resnet18, layer blocks: [4, 4, 4], channels: [8, 16, 32, 64], fuse mode: AsymBi
+UIUNet [18](TIP’2023) Channels: [64, 128, 256, 512], fuse mode: AsymBi
+MTUNet [41](TGRS’2023) Backbone: resnet18, layer blocks: [2, 2, 2, 2], filter: [16, 32, 64, 128, 256]
+SeRankDet [42](TGRS’2024) Channels:[64, 128, 256, 512, 1024],o=3
+SCTransNet [43](TGRS’2024) Backbone:Swin-Tiny, Layer Num=[2,6,6,2], Head Num=[3,6,12,24]
 $\star$ DCGANet (Ours) Channels: [64, 128, 256, 512, 1024]
+[…]
+IV Experiments
 IV-A Experimental Setup
 IV-A1 Datasets
-To rigorously evaluate the performance and generalization capabilities of our method, we conduct comprehensive benchmark experiments on three widely recognized infrared small target detection datasets: NUAA-SIRST [32], IRSTD-1K [31], and NUDT-SIRST [34].
-These datasets demonstrate considerable diversity in target characteristics, background clutter, and imaging conditions, thereby enabling robust validation of detection methodologies within real-world operational constraints.
+To rigorously evaluate the performance and generalization capabilities of our method, we conduct comprehensive benchmark experiments on three widely recognized infrared small target detection datasets: NUAA-SIRST, IRSTD-1K, and NUDT-SIRST.
+These datasets exhibit considerable diversity in target characteristics, background clutter, and imaging conditions, thereby enabling robust validation of detection methodologies.
+  • NUAA-SIRST: This dataset contains 427 real-world images capturing diverse scenes, such as clouds, cities, and oceans.
+With 480 target instances, it serves as a foundational benchmark for evaluating detection performance in common scenarios.
+  • IRSTD-1K: Comprising 1,001 highly diverse images, IRSTD-1K is designed to challenge a model’s cross-domain generalization.
+The complexity of its scene types and target shapes imposes significantly higher demands compared to other datasets.
+  • NUDT-SIRST: This synthetic dataset includes 1,327 images with complex backgrounds.
+By employing adaptive functions for target size and intensity blurring, it generates a wide variety of targets (point-like, patchy, extended), creating a distribution that closely mimics real-world data and tests model robustness against complex interference.
+IV-A2 Evaluation Metrics
+To quantitatively analyze our DCGANet and the comparative methods, we adopt four standard metrics: Intersection over Union (IoU), normalized IoU (nIoU), Probability of Detection ($P_{d}$), and False-alarm Rate ($F_{a}$).
+  • IoU measures the localization accuracy by comparing the overlap between the predicted bounding box and the ground truth. It is defined as the ratio of their intersection to their union.
+    $\text{IoU}=\frac{TP}{T+P-TP}$ (14)
+  • nIoU is a normalized adaptation of the IoU metric, averaging the IoU over all target-containing images to provide a dataset-level performance score.
+    $\text{nIoU}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP(i)}{T(i)+P(i)-TP(i)}$ (15)
+  • $P_{d}$ (also known as Recall) quantifies the model’s ability to correctly identify true targets. A higher value is better.
+    $P_{d}=\frac{TP}{TP+FN}$ (16)
+  • $F_{a}$ measures the rate of pixels incorrectly identified as targets. It is calculated per image, and a lower value is better.
+    $F_{a}=\frac{FP}{\text{Number of Pixels in Image}}$ (17)
+where $N$ is the total number of images with targets, $T$ and $P$ are the sets of ground-truth and predicted positive pixels, respectively.
+$TP$, $FP$, and $FN$ denote the counts of true positive, false positive, and false negative pixels.
+In addition, the Receiver Operating Characteristic (ROC) curve illustrates the trade-off between $P_{d}$ and $F_{a}$.
+IV-A3 Implementation Details
+We adopt U-Net as the baseline, Soft-IoU Loss as the loss function, and AdamW as the optimizer, incorporating a polynomial learning rate decay strategy.
+A deep supervision strategy is also employed to enhance performance.
+All experiments are conducted using distributed training on a system with four NVIDIA RTX 3090 GPUs.
+To ensure a fair and comprehensive evaluation, we train independent models for each of the three datasets.
+The hyperparameter configurations for all comparative methods are summarized in Table III-C.
+Due to significant differences in data distribution and resolution, we customize the training hyperparameters for each dataset, as detailed in Table II.
+TABLE II: Customized Training Hyperparameters for Each Dataset.
+[…]
+[…]
+IRSTD-1K 500 1e-4 4 $512\times 512$
+NUDT-SIRST 500 1e-4 16 $256\times 256$
+[…]
+IV-B Comparison with State-of-the-Art Methods
+We rigorously benchmark our DCGANet against a wide range of state-of-the-art (SOTA) methods, including traditional model-driven techniques (Background Suppression, HVS-based, and Low-rank Matrix-based) and various advanced deep learning models.
+The comprehensive results on all three datasets are summarized in Table IV-B, leading to several key observations:
   1. 1.
-    NUAA-SIRST: The real long-range infrared single-frame dataset comprises 427 images capturing diverse realistic scenes, such as clouds, cities, and oceans.
-However, the dataset is relatively small in scale, with only 480 target instances, which may limit its effectiveness for training and evaluating models in scenarios that demand high data diversity.
+    Superior Overall Performance: Our proposed DCGANet consistently achieves the best or second-best performance across all metrics on all three datasets.
+This demonstrates its superior accuracy and robustness in handling the challenges of low signal-to-noise ratio (SNR) and indistinct target boundaries.
+The combination of its dynamic feature extraction and fusion modules proves highly effective.
   2. 2.
-    IRSTD-1K: IRSTD-1K is designed to address the challenges of multi-scene cross-domain detection, comprising 1,001 highly diverse images.
-The complexity of its scene types and target shapes imposes significantly higher demands on the model’s generalization capabilities compared with the other two datasets.
+    Data-Driven vs.
+Model-Driven: A significant performance gap (often >15% in IoU) exists between deep learning (DL) methods and traditional approaches.
+This highlights the inherent limitations of model-driven techniques, which rely on fixed priors like local contrast or global sparsity.
+In contrast, data-driven DL methods show superior generalization by learning deep hierarchical representations, making them more adaptable and less sensitive to hyperparameters in complex, unstructured scenes.
   3. 3.
-    NUDT-SIRST: The scene-enhanced dataset, constructed from synthetic data, comprises 1,327 images with complex backgrounds, including cities, fields, bright regions, oceans, and clouds.
-By employing adaptive target size and intensity blurring functions, it generates diverse targets in the form of point-like, patchy, and extended shapes, yielding distribution characteristics that closely resemble those of real-world datasets.
-IV-A2 Evaluation Metrics
-To quantitative analyze our DCGANet and other methods, we adopt four standard metrics to evaluate detection performance: Intersection over Union (IoU), normalized IoU (nIoU), Probability of detection (Pd), and False alarm rate (Fa).
-The detailed description of these metrics are as follows:
-  1. 1.
-    IoU measures the accuracy of object localization by comparing the overlap between the predicted bounding box and the ground truth.
-It is defined as the ratio of the intersection area to the union area of the two bounding boxes.
-    $\text{IoU}=\frac{TP}{T+P-TP}$ (14)
-  2. 2.
-    nIoU is a standardized adaptation of the IoU metric.
-    $\text{nIoU}=\frac{1}{N}\sum_{i=1}^{N}\frac{TP(i)}{T(i)+P(i)-TP(i)}$ (15)
-  3. 3.
-    Pd quantifies the likelihood of correctly identifying true targets.
-    $\text{Pd}=\frac{TP}{TP+FN}$ (16)
-  4. 4.
-    Fa measures the frequency of incorrect target detections.
-    $\text{Fa}=\frac{FP}{FP+TN}$ (17)
-where $N$ is the total number of samples, $T$ and $P$ correspond to the number of ground truth and predicted positive pixels, respectively.
-Additionally, $TP$, $FP$, and $FN$ denote the counts of true positive, false positive, and false negative pixels, respectively.
-In addition, the ROC curve illustrates the variation trend of the Probability of detection (Pd) across different False alarm rate (Fa).
-IV-A3 Implementation Details
-We adopt U-Net as the baseline model, Soft-IoU Loss [53] as the loss function, and AdamW [54] as the optimizer, and incorporating a polynomial learning rate decay strategy to optimize the training process.
-Additionally, we employ a deep supervision strategy to further enhance model performance.
-All experiments are conducted using distributed training on a system equipped with four RTX 3090 GPUs.
-To ensure comprehensive and reliable evaluation, we train independent models on each of the three datasets.
-The detailed hyperparameter configurations for all comparative methods are summarized in Tab.
-I.
-To address the significant differences in dataset distribution and resolution, we customize the hyperparameters for each dataset, with specific configurations provided in Tab.
-II.
-TABLE II: Customized Hyperparameters for Each Dataset
-Dataset Epochs Lr Batch Resolution
-NUAA-SIRST 500 0.0001 4 512$\times$512
-IRSTD-1K 500 0.0001 4 512$\times$512
-NUDT-SIRST 500 0.0001 16 256$\times$256
-TABLE III: Comparison with Other State-of-the-art methods on four datasets.
-The $\uparrow$ indicates that the higher the indicator, the better, and the $\downarrow$ indicates that the lower the indicator, the better.
-The best performance is highlighted in bold, and the second-best is underlined.
-Method NUAA-SIRST(Tr=60%) IRSTD-1K(Tr=60%) NUDT-SIRST(Tr=60%)
-IoU $\uparrow$ nIoU $\uparrow$ $P_{d}$uparrow $F_{a}$downarrow IoU $\uparrow$ nIoU $\uparrow$ $P_{d}$uparrow $F_{a}$downarrow IoU $\uparrow$ nIoU $\uparrow$ $P_{d}$uparrow $F_{a}$downarrow
-                      Background Suppression Methods
-NWMTH [12] 15.35 17.00 71.10 55.21 18.31 16.74 49.97 21.06 18.87 16.22 62.84 32.45
-FKRW [47] 22.10 28.81 81.41 16.67 10.12 16.83 69.88 24.19 10.04 16.76 72.31 66.90
-HVS-based Methods
-MPCM [17] 25.64 27.63 68.01 45.07 20.72 23.26 58.78 31.10 27.40 38.95 55.54 11.82
-WLDM [48] 5.55 3.37 64.97 21.22 8.16 10.32 54.85 14.79 8.19 11.63 16.30 64.14
-RLCM [49] 25.56 27.46 86.34 62.16 16.46 21.83 70.86 67.70 8.21 11.42 70.22 31.67
-ILCM [50] 18.42 19.16 38.00 40.63 13.67 11.34 43.32 47.67 5.16 6.48 26.27 88.16
-TLLCM [41] 18.20 27.62 82.23 16.18 10.86 18.01 68.95 24.99 13.44 16.98 77.02 69.66
-GSWLCM [18] 15.67 13.82 72.48 21.96 12.81 13.46 70.57 13.82 21.08 19.92 69.06 42.78
-Low-rank and Sparse Decomposition
-IPI [42] 26.67 32.57 83.55 11.97 27.92 30.12 81.27 26.18 25.16 34.64 76.26 43.41
-RIPT [24] 11.85 14.91 79.48 22.67 14.71 17.43 79.55 28.35 24.13 33.98 78.54 56.24
-PSTNN [26] 22.72 23.50 77.36 29.99 28.83 29.09 82.68 39.48 21.26 24.22 72.78 61.26
-Deep Learning Methods
-ACM [32] 73.38 74.66 96.06 9.94 60.38 59.83 91.82 21.49 73.86 72.97 97.91 76.22
-AGPCNet [38] 77.86 75.23 98.92 34.95 68.30 66.18 94.29 15.55 88.55 88.08 97.40 34.09
-DNANet [34] 78.82 79.72 98.63 13.37 69.96 69.35 92.02 23.19 91.32 90.10 97.55 30.71
-ISNet[31] 80.83 78.55 98.78 4.93 68.74 61.98 95.80 14.97 88.02 88.76 93.14 28.24
-UIUNet [33] 79.76 78.74 98.69 4.41 65.20 64.57 92.33 13.45 88.86 88.11 98.40 23.17
-MTUNet [51] 78.41 79.09 98.02 8.66 67.04 68.24 93.18 12.86 87.39 87.34 98.15 29.43
-SCTransNet [52] 80.88 79.45 96.48 6.48 71.22 70.11 94.45 13.94 92.39 92.32 97.71 28.76
-SeRankDet [55] 81.26 79.66 98.74 2.77 74.45 70.03 95.12 10.65 94.16 93.34 98.71 22.36
-$\star$ DCGANet (Ours) 82.03 79.81 98.88 2.99 75.59 72.16 96.64 9.21 94.98 94.46 98.75 20.34
-IV-B Comparison With State-of-the-Arts
-Our DCGANet method is rigorously compared against state-of-the-arts methods across three datasets using metrics include IoU, nIoU, Pd, and Fa.
-The BS-based methods NWMTH [12], FKRW [47].
-The HVS-based Methods MPCM [17], WLDM [48], RLCM [49], ILCM [50], TLLCM [41], GSWLCM [18].
-The LRM-based methods IPI [42], RIPT [24], PSTNN [26].
-The deep learning methods ACM [32], AGPCNet [38], DNANet [34], ISNet [31], UIUNet [33], MTUNet [51], SCTransNet [52], SeRankDet [55].
-The experimental results summarized in Tab.
-III, including three critical observations:
-1) Our proposed DCGANet exhibits superior performance compared to state-of-the-art methods across multiple datasets.
-Specifically designed to tackle the challenges posed by low signal-to-noise ratio (SNR) and blurred target boundaries in infrared small target detection, our approach incorporates a multi-modal feature extractor that precisely extracts target features while effectively suppressing background clutter.
-Additionally, a dedicated feature fusion module is included to aggregate target details from various hierarchical levels.
-This innovative design substantially enhances detection accuracy and robustness, resulting in significant advancements in the field of infrared small target detection.
-2) The experimental results reveal a performance gap exceeding 15% in Intersection over Union (IoU) between Deep Learning-based methods and traditional approaches, highlighting the inherent limitations of model-driven techniques that rely on global sparsity or contrast priors.
-In contrast, data-driven DL methods demonstrate superior generalization capabilities by leveraging semantic feature extraction and deeper hierarchical representations.
-This allows them to significantly reduce sensitivity to hyperparameters in unstructured scenarios, resulting in more robust and adaptable performance.
-3) On the NUAA-SIRST dataset, DCGANet surpasses SeRankDet across all evaluation metrics except for the False alarm rate (Fa).
-While SeRankDet shows moderate effectiveness in combining global and local features, its performance significantly deteriorates in cluttered environments.
-In contrast, DCGANet addresses these limitations by leveraging precise target localization and dynamic feature fusion, resulting in superior pixel-level segmentation accuracy.
-Furthermore, Fig.
-6 illustrates receiver operating characteristic (ROC) curves for three datasets.
-The analysis corroborates the numerical results, confirming DCGANet’s stability and superiority over conventional techniques.
+    Advantage over SOTA DL Methods: On the challenging NUAA-SIRST and IRSTD-1K datasets, DCGANet surpasses recent powerful models like SeRankDet.
+For instance, on NUAA-SIRST, DCGANet achieves the highest IoU, nIoU, and $P_{d}$, with only a marginal trade-off in the $F_{a}$ rate compared to SeRankDet.
+This indicates that while other models may be effective, DCGANet’s architecture provides a better balance of high detection probability and precise pixel-level localization, especially in cluttered environments.
+Furthermore, Figure 6 illustrates the Receiver Operating Characteristic (ROC) curves for the three datasets.
+The analysis visually corroborates the quantitative results in Table IV-B, confirming that DCGANet consistently achieves a higher probability of detection ($P_{d}$) at any given false-alarm rate ($F_{a}$), underscoring its stability and superiority.
+TABLE III: Quantitative comparison with state-of-the-art methods on three datasets.
+For all metrics except $F_{a}$, higher is better ($\uparrow$).
+For $F_{a}$, lower is better ($\downarrow$).
+The best performance is highlighted in red, and the second-best is in blue.
+[…]
+[…]
+[…]
+[…]
+FKRW [37] 21.23 27.67 78.18 16.66 9.72 16.16 67.11 24.18 9.64 16.09 69.44 66.88
+[…]
+[…]
+WLDM [38] 5.33 3.23 62.40 21.21 7.84 9.91 52.68 14.79 7.87 11.17 15.65 64.11
+RLCM [39] 24.55 26.37 82.92 62.13 15.81 20.96 68.05 67.67 7.89 10.97 67.44 31.65
+ILCM [40] 17.69 18.40 36.50 40.61 13.13 10.89 41.60 47.65 4.96 6.22 25.23 88.20
+TLLCM [27] 17.48 26.53 78.98 16.17 10.43 17.30 66.22 24.98 12.91 16.31 73.97 69.63
+GSWLCM [12] 15.05 13.27 69.61 21.95 12.30 12.93 67.78 13.81 20.25 19.13 66.33 42.77
+[…]
+[…]
+RIPT [13] 11.38 14.32 76.33 22.66 14.13 16.74 76.40 28.34 23.18 32.63 75.43 56.21
+PSTNN [28] 21.82 22.57 74.29 30.00 27.69 27.94 79.41 39.46 20.41 23.26 69.89 61.24
+[…]
+[…]
+AGPCNet [22] 74.77 72.26 95.00 34.94 65.59 63.56 90.55 15.54 85.04 84.59 93.54 34.07
+DNANet [19] 75.70 76.57 94.73 13.37 67.19 66.60 88.38 23.18 87.70 86.53 93.69 30.69
+ISNet [32] 77.63 75.44 94.86 4.93 66.02 59.53 92.00 14.96 84.53 85.26 89.45 28.22
+UIUNet [18] 76.60 75.62 94.79 4.41 62.61 62.01 88.67 13.45 85.34 84.62 94.50 23.16
+MTUNet [41] 75.30 75.96 94.14 8.65 64.39 65.54 89.49 12.86 83.93 83.89 94.27 29.42
+SCTransNet [43] 77.67 76.30 92.66 6.48 68.39 67.34 90.71 13.94 88.73 88.66 93.84 28.75
+SeRankDet [42] 78.04 76.51 94.82 2.77 71.50 67.26 91.36 10.64 90.43 89.64 94.81 22.35
+[…]
+[…]
 (a) NUAA-SIRST
@@
 (c) NUDT-SIRST
-Figure 6: ROC curves of different methods on three datasets. Our DCGANet achieves the highest Pd at very low Fa.
+Figure 6: ROC curves of different methods on the three datasets.
+Our DCGANet consistently achieves the highest $P_{d}$ at very low $F_{a}$ levels, demonstrating its superior trade-off between detection rate and false alarms.
 IV-C Visualization Analysis
-To qualitatively validate the advantages of DCGANet, we analyze representative samples from the three datasets (Fig.
-7 - 8).
-Traditional model-driven methods, which rely on handcrafted features, exhibit high false alarms rates (Fig.
-7, yellow boxes) and frequent missed detections (Fig.
-8, cyan boxes).
-In contrast, deep learning-based methods leverage learned hierarchical representations to distinguish targets from complex background clutter effectively, showcasing superior detection robustness.
-Gradient-weighted class activation mapping (GradCAM) visualizations for DCGANet (Fig.
-9) highlight the focus for each module during the target extraction.
-The Selective Variable Convolution (SVC) employs diverse convolution forms compared to the baseline UNet, enabling robust feature preservation and edge refinement.
-The dynamic content guided attention (DCGA) module adaptively refines the distinction between target and background, demonstrating enhanced precision in focusing on true targets compared to static attention mechanisms.
-Additionally, the Adaptive Dynamic Feature Fusion (ADFF) module improves edge preservation, achieving sub-pixel accuracy in localization.
-These results collectively demonstrate that DCGANet’s modular, with each component exhibiting clear and discriminative attention patterns.
-Figure 7: Visualization comparison of detection results via different methods on representative images from NUAA-SIRST, IRSTD-1K, and NUDT-SIRST datasets, indicate the land forests, oceans skys, and city building interfere.
-The red, green, and yellow boxes denote correct detection, false alarms, and missed detections, respectively.
-Figure 8: Visualization comparesion the detection results of different methods (The red, green, and yellow boxes denote correct detection, false alarms, and missed detections).
-3D mesh plot highlighting the features of small targets in the image.
+To qualitatively validate the advantages of DCGANet, we analyze representative samples from the three datasets (Figures 7 and 8).
+Traditional model-driven methods, which rely on handcrafted features, exhibit high false alarm rates (yellow boxes) and frequent missed detections.
+In contrast, deep learning-based methods leverage learned hierarchical representations to distinguish targets from complex background clutter more effectively, showcasing superior detection robustness.
+Gradient-weighted Class Activation Mapping (GradCAM) visualizations for DCGANet (Figure 9) highlight the focus area for each module during target extraction.
+The SVC module employs diverse convolutional forms, enabling robust feature preservation and edge refinement compared to the baseline.
+The DCGA module adaptively refines the distinction between target and background, demonstrating enhanced precision in focusing on true targets.
+Finally, the ADFF module improves edge preservation by dynamically fusing features, achieving sub-pixel accuracy in localization.
+These results collectively demonstrate that DCGANet’s modules operate synergistically, with each component exhibiting clear and discriminative attention patterns.
+Figure 7: Visualization comparison of detection results on representative images from the three datasets, showcasing interference from land, oceans, and cityscapes.
+The red, green, and yellow boxes denote correct detections, false alarms, and missed detections, respectively.
+Figure 8: Further visualization comparisons.
+The main images show detection results, while the 3D mesh plots below each one highlight the intensity profile of the small target feature against its local background.
+Figure 9: Qualitative comparison of attention maps on infrared images.
+From left to right: original images, predictions by U-Net (baseline), U-Net+SVC, U-Net+SVC+DCGA, U-Net+SVC+DCGA+ADFF (our full DCGANet), and ground-truth masks.
+The improvements in target localization and detail suppression of noise are evident with the addition of each module.
 IV-D Ablation Study
-IV-D1 Model-wise performance
-Ablation experiments were conducted to evaluate the proposed SVC, DCGA, and ADFF modules, demostrating their individual and integrated impacts on model performance, as shown in Tab.
-IV.
-The baseline U-Net was used for comparison analysis.
+IV-D1 Component-wise Performance Analysis
+Ablation experiments were conducted to systematically evaluate the proposed SVC, DCGA, and ADFF modules, demonstrating their individual and integrated impacts on model performance.
+As shown in Table IV, the baseline U-Net was used for comparison.
   1. 1.
-    Independent efficacy: Each module yields distinct performance improvements, corroborating its theoretical motivations (e.g. SVC’s 5.50% IoU gain on IRSTD-1K and DCGA’s 7.83% improvement).
+    Independent Efficacy: Each module individually yields distinct performance improvements over the baseline (a), corroborating its theoretical motivation.
+For instance, adding DCGA alone (c) improves IoU by 7.67% on IRSTD-1K.
   2. 2.
-    Convolutional Advancement: Replacing vanilla convolutions with the SVC module elevates IoU by 4.62% on NUAA-SIRST and 5.50% on IRSTD-1K, demonstrating its superiority feature encoding multi-scale target features and suppressing background clutter.
+    Effectiveness of SVC: Replacing standard convolutions with our SVC module (b) elevates the IoU by 4.53% on NUAA-SIRST and 7.35% on IRSTD-1K, proving its superiority in encoding multi-scale target features while suppressing clutter.
   3. 3.
-    Attention Mechanism Superiority:The DCGA outperforms established attention modules (e.g.
-SE, DAM, and CBAM) by adaptively modeling channel-specific spatial importance maps (SIMs) critical for sub-pixel infrared targets.
-  4. 4.
-    Contextual Fusion:The ADFF module’s dynamic feature fusion strategy, enhances detection accuracy, achieving a 9.62% improvement over baseline fusion strategies on the IRSTD-1K dataset.
-This demonstrates the effectiveness of leveraging broader contextual information for target detection in complex infrared scenes.
-Furthermore, Fig.
-9 presents a qualitative comparison of detection results under different model configurations.
-The visualizations include original infrared inputs, ground-truth masks, attention maps from the baseline U-Net, and results from the progressively integrated DCGANet, showcasing the incremental benefits of each module in target localization and feature enhancement.
-Figure 9: Qualitative comparison of detection results on infrared images.
-From left to right: original images, predictions by U-Net, U-Net+SVC, U-Net+SVC+DCGA, U-Net+SVC+DCGA+ADFF/DCGANet, and ground-truth masks.
-The improvements in target localization and detail are evident with the addition of each module.
-TABLE IV: Quantitative Ablation Study on DCGANet (selective variable convolution is marked SVC, dynamic content guided attention is marked DCGA, adaptive dynamic feature fusion module is marked ADFF), The $\uparrow$ indicates that higher values are better.
-Strategy Module NUAA-SIRST IRSTD-1K NUDT-SIRST
-SVC DCGA ADFF IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$
-  (a) 75.12 75.35 64.42 62.54 89.81 89.58
-(b) 79.74 78.53 71.92 69.51 92.42 92.39
-(c) 80.11 78.75 72.25 70.21 93.02 92.88
-(d) 81.34 79.01 74.86 72.42 93.98 93.52
-(e) $\checkmark$ $\checkmark$ $\checkmark$ 82.03 79.81 75.59 72.16 94.98 94.46
-IV-D2 Ablation of SVC module
-The SVC module integrates deformable convolutio (DC) and multi-dilation convolution (MDC) alongside Standard Convolution to generate more robust feature representations.
-Ablation results in Tab.
-V reveal that while DC alone provides a moderate performance gain, the integration of MDC yields significantly greater improvementds.
-This demonstrates the superior efficacy of the SVC composite convolutional block, where multi-rate dilated convolutions play a pivotal role in expanding receptive fields and capturing hierarchical contextual information.
-These findings underscore the critical importance of broad receptive fields in IRSTD, serving as a key motivator for the design of the ADFF module.
-By leveraging broader spatial context, the ADFF module is engineered to refine target detection further, enhancing both accuracy and robustness in complex infrared scenes.
-This synergy between the SVC’s multi-convolutional feature extraction and the ADFF’s contextual fusion highlights the importance of hierarchical receptive field design in overcoming the challenges of low SNR and ambiguous target boundaries.
-TABLE V: Ablation experiments of convolution branches of SVC.
-Strategy Module NUAA-SIRST IRSTD1K NUDT-SIRST
-SC MDC DC IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$
-(a) 80.35 77.59 72.14 70.12 91.21 91.96
-(b) 81.72 78.47 74.01 70.85 93.30 93.88
-(c) 80.97 78.22 73.45 70.36 92.64 93.37
-(d) 82.03 79.81 75.59 72.16 94.98 94.46
-IV-D3 Ablation of DCGA module
-Traditional attention modules (SE/DAM/CBAM) fail to address the challenges specific to IRSTD:
-  1. 1.
-    Under the condition of low SNR, single-channel simulation cannot adequately represent cross-channel variance.
-  2. 2.
-    Compared with the serial design, the channel shuffling architecture of parallel processing DCGA improves the small target recognition rate.
-As shown in Tab.
-VI, in detection performance, DCGA has significantly improved compared with SE, DAM and CBAM.
-This shows that DCGA is able to better recalibrate the feature distribution by learning Spatial Importance Maps (SIM) to focus on signal differences between channel-wise, which is particularly critical for enhancing infrared small target features.
-TABLE VI: Ablation of DCGA compared with other mainstream attention mechanisms
-Strategy NUAA-SIRST IRSTD1K NUDT-SIRST
-IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$ IoU $\uparrow$ nIoU $\uparrow$
-SE 78.74 77.90 68.13 68.47 90.93 90.98
-DAM 79.88 78.01 70.19 69.11 91.29 91.65
-CBAM 80.00 77.88 71.20 68.59 92.30 92.04
-DCGA 82.03 79.81 75.59 72.16 94.98 94.46
-TABLE VII: Ablation study of DCGA. All the experiments are conducted on three datasets.
-Method DCGAcas DCGAw/o r DCGAw/o CS DCGA
-parallel CA&SA
-cascaded CA&SA
-refine ✓(remove CS)
-NUAA-SIRST IoU 80.12 75.15 78.54 82.03
-nIoU 78.71 75.24 77.12 79.81
-IRSTD-1K IoU 72.25 64.41 71.34 75.59
-nIoU 70.20 62.97 69.04 72.16
-NUDT-SIRST IoU 93.04 89.14 91.19 94.48
-nIoU 92.78 89.43 91.01 94.46
-Further ablation studies on the DCGA design settings are as follows:
-  1. 1.
-    DCGAcas represents the cascaded connection of channel attention (CA) and spatial attention (SA) in contrast to the parallel mode of DCGA;
-  2. 2.
-    DCGAw/o r denotes the parallel connection of channel attention (CA) and spatial attention (SA) but without the refinement process;
-  3. 3.
-    DCGAw/o CS indicates the parallel connection of channel attention (CA) and spatial attention (SA) but excludes the Channel Shuffle (CS) operation in the refinement stage;
-  4. 4.
-    DCGA is the proposed method that integrates parallel connection of channel attention (CA) and spatial attention (SA), along with a complete refinement process that includes channel shuffle (CS).
-The results are presented in Tab.
-VII, which reveals that model performance slightly declines when the channel shuffle (CS) operation is removed.
-This result indicates that CS plays a crucial role in facilitating cross-channel information interaction to highlight small target features.
-Additionally, the channel attention (CA) and spatial attention (SA) mechanisms in DCGA are deployed in parallel, efficiently modeling the salient regions of small targets by generating initial spatial importance maps (SIMs) in a parallel manner.
-We also experimented with a cascaded design, and the results show that both parallel and cascaded designs exhibit very similar performance in terms of signal-to-noise ratio (SNR) gain.
-However, when the refinement process is omitted (DCGAw/o CS), performance significantly deteriorates, further confirming the importance of the refinement stage in enhancing the feature representation of infrared small targets.
+    Synergy of Modules: Combining SVC and DCGA (d) provides a significant boost over using either module alone.
+The full model, DCGANet (e), which also includes ADFF, achieves the overall best performance, demonstrating the powerful synergy of the three proposed components.
+Furthermore, Figure 9 presents a qualitative comparison of attention maps under different model configurations.
+The visualizations showcase the incremental benefits of each module in refining target localization and enhancing feature representation.
+TABLE IV: Quantitative ablation study of DCGANet’s core components.
+We measure the performance improvement by sequentially adding Selective Variable Convolution (SVC), Dynamic Context-Guided Attention (DCGA), and Adaptive Dynamic Feature Fusion (ADFF) to a U-Net baseline.
+[…]
+[…]
+[…]
+(b) 76.59 75.42 69.07 66.76 88.76 88.73
+(c) 76.94 75.64 69.39 67.43 89.34 89.20
+(d) 78.12 75.88 71.89 69.55 90.26 89.82
+(e) 78.78 76.65 72.60 69.31 91.22 90.72
+[…]
+IV-D2 Ablation of SVC Module Design
+The SVC module integrates Standard Convolution (SConv), Deformable Convolution (DConv), and Multiple Dilation Convolution (MDConv).
+To analyze each branch’s contribution, we conduct an ablation study shown in Table V.
+The results reveal that while adding DConv provides a moderate performance gain (c), the addition of MDConv yields significantly greater improvements (b).
+The full three-branch SVC (d) achieves the best performance, demonstrating that multi-rate dilated convolutions play a pivotal role in expanding the receptive field and capturing crucial hierarchical context.
+TABLE V: Ablation study of the convolutional branches within the SVC module.
+We start with a standard convolution (SConv) and add the other branches.
+Results on NUDT-SIRST are included for comprehensive analysis.
+[…]
+[…]
+[…]
+(b) 78.49 75.36 71.08 68.04 90.46 90.31
+(c) 77.76 75.13 70.54 67.57 89.25 89.03
+(d) 78.78 76.65 72.60 69.31 91.22 90.72
+[…]
+IV-D3 Ablation of DCGA Module Design
+We compare DCGA against other mainstream attention mechanisms (SE, DAM, CBAM) in Table VI.
+DCGA significantly outperforms them, showing its ability to better recalibrate feature distributions by learning channel-specific Spatial Importance Maps (SIMs), which is critical for enhancing infrared small target features.
+TABLE VI: Comparison of DCGA with other mainstream attention mechanisms.
+[…]
+[…]
+[…]
+DAM 76.71 74.92 67.41 66.38 87.67 88.02
+CBAM 76.83 74.79 68.38 65.88 88.64 88.40
+DCGA (Ours) 78.78 76.65 72.60 69.31 91.22 90.72
+[…]
+Furthermore, we conducted a detailed ablation of the DCGA’s internal design choices (Table VII). The study validates our design:
+  • DCGAcas: A cascaded connection of Channel (CA) and Spatial Attention (SA) performs slightly worse than our parallel design, suggesting parallel processing is more effective for this task.
+  • DCGAw/o r: Removing the refinement stage entirely causes a significant performance drop, confirming the importance of the refinement process.
+  • DCGAw/o CS: Removing only the Channel Shuffle (CS) from the refinement stage also degrades performance, indicating that CS is crucial for cross-channel information interaction.
+TABLE VII: Detailed ablation study on the internal design of the DCGA module.
+We compare our full design against three variants: a cascaded version (DCGAcas), one without the refinement stage (DCGAw/o r), and one without the channel shuffle operation (DCGAw/o CS).
+The best results are highlighted in bold.
+[…]
+[…]
+nIoU (%) $\uparrow$ 75.60 72.26 74.07 76.65
+[…]
+nIoU (%) $\uparrow$ 67.42 60.48 66.31 69.31
+[…]
+nIoU (%) $\uparrow$ 89.10 85.89 87.41 90.72
+[…]
 V Conclusion
-In this paper, we propose a dynamic content guided attention multiscale feature aggregation network (DCGANet), which adopts a “coarse-to-fine” attention principle to achieve high detection accuracy.
+In this paper, we propose a dynamic content-guided attention multiscale feature aggregation network (DCGANet), which adopts a "coarse-to-fine" attention principle to achieve high detection accuracy.
 First, the selective variable convolution (SVC) module is introduced, integrating standard convolution, deformable convolution, and multi-rate dilated convolution.
 This module expands the receptive field and enhances non-local features, thereby effectively improving the discrimination between targets and backgrounds.
-Second, the core component of DCGANet is the dynamic content guided attention (DCGA) module, which uses two-stage attention to first direct the network’s focus to significant regions in feature maps and then determine whether these regions correspond to targets or background interference.
+Second, the core component of DCGANet is the Dynamic Content-Guided Attention (DCGA) module, which uses two-stage attention to first direct the network’s focus to significant regions in feature maps and then determine whether these regions correspond to targets or background interference.
 By retaining the most salient responses, this mechanism suppresses false alarms effectively.